### v3

# Zagreb, Croatia Ensembling uncertainty with negative objectness

Abstract
Outlier detection is an essential capability of safety-critical visual recognition. Many existing methods deliver good results by encouraging standard closed-set models to produce low-confidence predictions in negative training data. However, that approach conflates prediction uncertainty with recognition of outliers. We disentangle the two factors by revisiting the K+1-way classifier that involves K known classes and one negative class. This setup allows us to formulate a novel outlier score as an ensemble of in-distribution uncertainty and the posterior of the negative class that we term negative objectness. Our UNO score can detect outliers due to either high prediction uncertainty or similarity with negative training data. We showcase the utility of our method in experimental setups with K+1-way image classification and K+2-way dense prediction. In both cases we show that the bias of real negative data can be relaxed by leveraging a jointly trained normalizing flow. Our models outperform the current state-of-the art on standard benchmarks for image-wide and pixel-level outlier detection.
1 Introduction
Discriminative models may become robust to outliers through upgrade to open-set recognition [Liang et al.(2022)Liang, Wang, Miao, and Yang]. This can be conveniently achieved by complementing standard classification with outlier detection [Zhang et al.(2020)Zhang, Li, Guo, and Guo]. Typically, the outlier detector delivers a scalar outlier score that induces ranking and enables detection through thresholding [Ruff et al.(2021)Ruff, Kauffmann, Vandermeulen, Montavon, Samek, Kloft, Dietterich, and Müller]. Many outlier detectors both in the image-wide [Hendrycks et al.(2018)Hendrycks, Mazeika, and Dietterich, Dhamija et al.(2018)Dhamija, Günther, and Boult] and the pixel-level context [Bevandić et al.(2022)Bevandić, Krešo, Oršić, and Šegvić, Biase et al.(2021)Biase, Blum, Siegwart, and Cadena, Chan et al.(2021b)Chan, Rottmann, and Gottschalk, Grcic et al.(2022)Grcic, Bevandic, and Segvic, Tian et al.(2022)Tian, Liu, Pang, Liu, Chen, and Carneiro] rely on negative training data. Although it cannot represent the entire variety of the visual world, the negative data can still help by signaling that not all data should be confidently recognized. If the variety of the negative data exceeds the variety of the inliers, then there is a reasonable hope that the test outliers will be detected. Unfortunately, this entails undesired bias towards test outliers that appear similar to the negative training data. This concern can be addressed either by relying on synthetic negatives [Lee et al.(2018)Lee, Lee, Lee, and Shin, Neal et al.(2018)Neal, Olson, Fern, Wong, and Li], or through separate ranking with respect to approaches that do not train on real negative data [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al., Blum et al.(2021)Blum, Sarlin, Nieto, Siegwart, and Cadena, Chan et al.(2021a)Chan, Lis, Uhlemeyer, Blum, Honari, Siegwart, Fua, Salzmann, and Rottmann].
Figure 1: We propose UNO, a plug-in module that enables outlier-aware inference atop a desired feature extractor. UNO decouples in-distribution uncertainty from outlier recognition, and boosts the outlier detection performance by ensembling the two components.
Many existing methods improve open-set performance of existing closed-set models by encouraging low-confidence predictions in negative training data [Hendrycks et al.(2022)Hendrycks, Basart, Mazeika, Zou, Kwon, Mostajabi, Steinhardt, and Song, Lee et al.(2018)Lee, Lee, Lee, and Shin]. However, this approach conflates prediction uncertainty with outlier recognition. In this paper, we disentangle the two factors by reconsidering the K+1-way classifier where the additional class represents the training negative data. This allows us to formulate the outlier score as the posterior of the K+1-th class, which we term negative objectness. Moreover, the outlier score can also be formulated as prediction uncertainty over the K known classes [Hendrycks et al.(2022)Hendrycks, Basart, Mazeika, Zou, Kwon, Mostajabi, Steinhardt, and Song]. We find that these two formulations exhibit complementary behaviour due to being sensitive to different test outliers. Consequently, we propose a novel outlier score that we term UNO: ensemble of Uncertainty over K inlier classes [Hendrycks and Gimpel(2017)] and Negative Objectness.
Our UNO score is a strikingly good fit in the pixel-level context as an extension of the direct set-prediction approach [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar, Yu et al.(2022)Yu, Wang, Qiao, Collins, Zhu, Adam, Yuille, and Chen, Li et al.(2023)Li, Zhang, xu, Liu, Zhang, Ni, and Shum]. We express the negative objectness as the posterior of the negative class in the frame of a K+2-way mask-level classifier. A closer look suggests that our score performs well due to different inductive bias of the two components as indicated by a remarkably weak correlation. The UNO score performs competitively in the image-wide context as well, even though the absence of the no-object class decreases the synergy of the two components. Our models outperform the state-of-the-art on several standard benchmarks in both learning setups, i.e. with and without real negative data.
2 Related work
Image-wide classification. Image-wide classification models can be complemented with outlier detection either through a modified training procedure or in a post-hoc manner with frozen parameters [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al.]. Post-hoc methods build upon pre-trained closed-set classifiers. Early baselines express the outlier score as maximum softmax probability [Hendrycks and Gimpel(2016)] or maximum logit [Hendrycks et al.(2019a)Hendrycks, Basart, Mazeika, Zou, Kwon, Mostajabi, Steinhardt, and Song]. TempScale [Guo et al.(2017)Guo, Pleiss, Sun, and Weinberger] calibrates softmax probabilities with temperature scaling while ODIN [Liang et al.(2017)Liang, Li, and Srikant] pre-processes the input with anti-adversarial perturbations. Recent post-hoc methods simplify layer activations [Djurisic et al.(2022)Djurisic, Bozanic, Ashok, and Liu] or consider distances from class prototypes [Zhang et al.(2022)Zhang, Fu, Chen, Du, Li, Wang, Han, Zhang, et al.]. In contrast, training methods involve various regularizations that either consider only inlier data or utilize real negative data as well. The former rely on modeling confidence [DeVries and Taylor(2018)], self-supervised training [Hendrycks et al.(2019b)Hendrycks, Mazeika, Kadavath, and Song, Tack et al.(2020)Tack, Mo, Jeong, and Shin], per-sample temperature scaling [Hsu et al.(2020)Hsu, Shen, Jin, and Kira], outlier synthesis in feature space [Du et al.(2022)Du, Wang, Cai, and Li, Tao et al.(2023)Tao, Du, Zhu, and Li] or mitigating overconfidence [Wei et al.(2022)Wei, Xie, Cheng, Feng, An, and Li]. Recent methods guide the learning process with respect to nearest neighbours [Park et al.(2023)Park, Jung, and Teoh] or leverage distances in feature space either by combining different distance functions [Noh et al.(2023)Noh, Jeong, and Lee] or distances in different feature spaces [Chen et al.(2023)Chen, Lin, Xu, and Vela]. Early work in training with negative data [Hendrycks et al.(2018)Hendrycks, Mazeika, and Dietterich, Dhamija et al.(2018)Dhamija, Günther, and Boult] encourages classifiers to produce high entropy in inlier samples. MCD [Yu and Aizawa(2019)] promotes discrepancy between prediction on OOD samples of two classification heads. UDG [Yang et al.(2021)Yang, Wang, Feng, Yan, Zheng, Zhang, and Liu] groups negative data following inlier taxonomy to enrich semantic knowledge of inlier classes. MixOE [Zhang et al.(2023a)Zhang, Inkawhich, Linderman, Chen, and Li] interpolates between inliers and negatives to improve regularization. Unlike related training based methods UNO does not modify the standard classifier training objective, but merely requires a slight architectural change to extend the classifier with a logit for the negative class. Additionally, UNO requires significantly less negative data than related methods.
Dense prediction. Many image-wide outlier detection methods can directly complement models for pixel-level prediction [Hendrycks and Gimpel(2016), Hendrycks et al.(2019a)Hendrycks, Basart, Mazeika, Zou, Kwon, Mostajabi, Steinhardt, and Song]. Early approaches estimate the prediction uncertainty with maximum softmax probability [Hendrycks and Gimpel(2017)], ensembling [Lakshminarayanan et al.(2017)Lakshminarayanan, Pritzel, and Blundell] or Bayesian uncertainty [Mukhoti and Gal(2018)]. Subsequent work suggests using energy scores [Liu et al.(2020)Liu, Wang, Owens, and Li] or standardized max-logit [Jung et al.(2021)Jung, Lee, Gwak, Choi, and Choo]. Training with negative data can be incorporated by pasting negative content from broad negative datasets (usually ADE20K [Zhou et al.(2017)Zhou, Zhao, Puig, Fidler, Barriuso, and Torralba] or COCO [Lin et al.(2014a)Lin, Maire, Belongie, Hays, Perona, Ramanan, Dollár, and Zitnick]) atop inlier images [Biase et al.(2021)Biase, Blum, Siegwart, and Cadena, Chan et al.(2021b)Chan, Rottmann, and Gottschalk, Grcic et al.(2022)Grcic, Bevandic, and Segvic, Hendrycks et al.(2018)Hendrycks, Mazeika, and Dietterich, Tian et al.(2022)Tian, Liu, Pang, Liu, Chen, and Carneiro, Bevandić et al.(2022)Bevandić, Krešo, Oršić, and Šegvić]. However, real negative data can be replaced with synthetic negative patches obtained by sampling a jointly trained generative model [Lee et al.(2018)Lee, Lee, Lee, and Shin, Neal et al.(2018)Neal, Olson, Fern, Wong, and Li, Zhao et al.(2021)Zhao, Cao, and Lin]. Outlier detection performance can be further improved by providing more capacity [Vaze et al.(2021)Vaze, Han, Vedaldi, and Zisserman]. Another branch of outlier detection methods considers generative models [Blum et al.(2021)Blum, Sarlin, Nieto, Siegwart, and Cadena, Grcic et al.(2022)Grcic, Bevandic, and Segvic, Liang et al.(2022)Liang, Wang, Miao, and Yang].
Mask-level recognition Recent panoptic architectures [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar, Yu et al.(2022)Yu, Wang, Qiao, Collins, Zhu, Adam, Yuille, and Chen, Li et al.(2023)Li, Zhang, xu, Liu, Zhang, Ni, and Shum] decompose scene understanding into class-agnostic segmentation and region-wide recognition. They frame the detection of semantic masks as direct set-prediction where each mask is classified into K inlier classes and one no-object class. This induces a degree of openness and therefore favours outlier detection. Recent work shows great applicability of the Mask2Former architecture for open-set segmentation [Nayal et al.(2023)Nayal, Yavuz, Henriques, and Güney, Rai et al.(2023)Rai, Cermelli, Fontanel, Masone, and Caputo, Ackermann et al.(2023)Ackermann, Sakaridis, and Yu, Grcić et al.(2023)Grcić, Šarić, and Šegvić]. All these methods detect anomalies due to not belonging to any inlier class and improve when training with negative data. Specifically, RbA [Nayal et al.(2023)Nayal, Yavuz, Henriques, and Güney] reduces energy in negative pixels, Mask2Anomaly [Rai et al.(2023)Rai, Cermelli, Fontanel, Masone, and Caputo] uses a contrastive loss, while EAM [Grcić et al.(2023)Grcić, Šarić, and Šegvić] ensembles region-wide outlier scores. However, these approaches conflate negative data with no-object regions. Furthermore, they fail to exploit the fact that mask queries behave like one-vs-all classifiers [Nayal et al.(2023)Nayal, Yavuz, Henriques, and Güney, Ackermann et al.(2023)Ackermann, Sakaridis, and Yu], which makes them appropriate for direct detection of anomalous regions. UNO addresses these weaknesses by decoupling negatives from the no-object class, and representing them as a standalone class in the K+2-way taxonomy.
3 Disentangling negative objectness from uncertainty
Let $h_{\theta_{1}}:\mathcal{X}\rightarrow\mathbb{R}^{d}$ be a feature extractor (e.g. ResNet [He et al.(2016)He, Zhang, Ren, and Sun] or ViT [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby]) onto which we attach our K+1-way classifier $g_{\theta_{2}}:\mathbb{R}^{d}\rightarrow\Delta^{K}$, where $\Delta^{K}$ is a K-dimensional probabilistic simplex. Given an input $\mathbf{x}\in\mathcal{X}$, we compute a latent representation $\mathbf{z}=h_{\theta_{1}}(\mathbf{x})$ and map it onto the simplex point $\mathbf{p}=\text{softmax}(g_{\theta_{2}}(\mathbf{z}))$. Thus, our deep model $f_{\theta}=g_{\theta_{2}}\circ h_{\theta_{1}}$ maps input samples from $\mathcal{X}$ onto simplex points in $\Delta^{K}$. We assume that the classifier $g_{\theta_{2}}$ is a simple projection. Thus we have $g_{\theta_{2}}(\textbf{x})=\textbf{W}\cdot\textbf{x}$, where the rows of W correspond to class vectors $\textbf{w}_{j}$.
Negative objectness. Given feature representation $\mathbf{z}$ and a linear classifier $g$, we formulate our negative objectness score ($\textbf{s}_{\text{NO}}$) as posterior of the K+1-th class:
$\mathbf{s}_{\text{NO}}(\mathbf{z}):=P(Y=\text{K+1}|\mathbf{z})=\frac{\text{exp}(\mathbf{w}_{K+1}^{T}\mathbf{z})}{\sum_{j=1}^{K+1}\text{exp}(\mathbf{w}_{j}^{T}\mathbf{z})}\;.$ (1)
The $\textbf{s}_{\text{NO}}$ score tests whether a given sample is semantically dissimilar from the inlier classes and similar to the negative training data. Consequently, a test outlier that is similar to the training negatives will yield high $\textbf{s}_{\text{NO}}$.
Prediction uncertainty. Negative objectness is not suitable for detecting test outliers that are semantically dissimilar from the training data. However, many of these outliers yield uncertain predictions across inlier classes of our K+1-classifier. We address this observation by defining the uncertainty score as negative prediction confidence across the K inlier classes:
$\mathbf{s}_{\text{Unc}}(\mathbf{z}):=-\max_{k=1...K}P(Y=k|\mathbf{z})=-\max_{k=1...K}\frac{\text{exp}(\mathbf{w}_{k}\mathbf{z})}{\sum_{j=1}^{K+1}\text{exp}(\mathbf{w}_{j}\mathbf{z})}=-\frac{\text{exp}(\mathbf{w}_{\hat{k}}\mathbf{z})}{\sum_{j=1}^{K+1}\text{exp}(\mathbf{w}_{j}\mathbf{z})}\;.$ (2)
The last equality in (2) denotes the index of the winning class as $\hat{k}=\text{argmax}_{k}\,\exp(\mathbf{w}_{k}^{T}\mathbf{z})$. The uncertainty score $\textbf{s}_{\text{Unc}}$ detects different outliers than negative objectness $\textbf{s}_{\text{NO}}$ despite attending to the same shared features $\mathbf{z}$.
UNO score. We integrate the prediction uncertainty (2) and negative objectness (1) into the UNO score as the sum of the two terms:
$\mathbf{s}_{\text{UNO}}(\mathbf{z})=\mathbf{s}_{\text{Unc}}(\mathbf{z})+\mathbf{s}_{\text{NO}}(\mathbf{z}).$ (3)
UNO has an interesting geometrical interpretation in the pre-logit space $\mathbb{R}^{d}$. We note that class vectors tend to be mutually orthogonal: $\mathbf{w}_{i}^{T}\mathbf{w}_{j}=0,\,\,\forall i,j\in\{1,2,...,K+1\}$. This arrangement is enforced by the standard supervised loss ($-\log\text{softmax}$) whenever there is enough data, as well as enough capacity and dimensionality ($d>K+1$) in the feature extractor $h_{\theta_{1}}$. Thus, a small angle between the latent representation z and the negative vector $\textbf{w}_{K+1}$ leads to high outlier objectness $\textbf{s}_{\text{NO}}$. Moreover, we note that the uncertainty score $\textbf{s}_{\text{Unc}}$ is negatively correlated with the feature norm $\lvert\lvert\mathbf{z}\rvert\rvert$. If we fix the direction of z, its norm can be viewed as reciprocal softmax temperature. As usual, large temperatures lead to uncertain predictions [Hinton et al.(2015)Hinton, Vinyals, and Dean].
The above insights show that the two components of our score will be at least partially decorrelated due to being sensitive to different outliers. Thus, the UNO score is likely to achieve lower error than any of the two components alone. Figure 2 illustrates our insight in an image-wide recognition experiment on small images. Further empirical analysis indicates a weak positive correlation between the UNO components, as detailed in the Appendix.
Figure 2: Interpretation of UNO in the pre-logit space $\mathbb{R}^{d}$ in an OpenOOD [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al.] experiment with CIFAR-10 inliers. Small $L_{2}$ norm of feature representations $\mathbf{z}$ leads to high $\textbf{s}_{\text{Unc}}$ while the angle between $\mathbf{z}$ and the K+1-th weight vector $\mathbf{w}_{\text{K+1}}$ leads to high $\textbf{s}_{\text{NO}}$. The two components capture different outliers as shown on the rightmost plot.
Extending UNO for dense prediction. The UNO outlier detector can be attached to any pre-trained deep classifier. In the case of image-wide prediction, we simply extend a pre-trained K-way classifier with an additional class. However, in the case of dense prediction, many contemporary architectures rely on mask-level recognition [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar]. Consequently, we build our per-pixel score by merging multiple mask-level UNO scores:
$\textbf{s}_{\text{UNO}}^{\text{M2F}}=\sum_{i=1}^{N}\mathbf{m}_{i}\cdot\textbf{s}_{\text{UNO}}(\textbf{z}_{i})\;.$ (4)
Equation 4 assumes that a dense prediction model has partitioned the input image into $N$ different masks $\mathbf{m}_{i}$ with the corresponding latent mask embeddings $\mathbf{z}_{i}$.
We note that the mask-level classifiers already use the K+1-th logit to indicate that the particular query has not been associated with any image region. It would be a bad idea to reuse that logit for mask-level negative objectness since unused queries are regular occurrences in in-distribution images. Consequently, we extend the default classifier with the negative class. In our implementations, the K+1-th logit corresponds to negative objectness while the K+2-th logit indicates the no-object class. Please find more details in the Appendix.
4 Training UNO with and without real negative data
Let $\mathcal{D}_{\text{in}}=\{(\textbf{x}_{i},y_{i})\}_{i=1}^{N^{+}}$ be an inlier dataset with $\mathbf{x}\in\mathcal{X}$ and $y_{i}\in\mathcal{Y}$, where $\mathcal{Y}$ is the set of classes of size $K=|\mathcal{Y}|$. Let $\mathcal{D}_{\text{out}}=\{\textbf{x}_{j}\}_{j=1}^{N^{-}}$ be a negative dataset that mimics test outliers. We construct a mixed-content training set $\mathcal{D}$ as the union of inlier samples with negative samples labeled as the K+1-th class: $\mathcal{D}=\mathcal{D}_{\text{in}}\cup\{(\mathbf{x}_{j},K+1)\,|\,\textbf{x}_{j}\in\mathcal{D}_{\text{out}}\}$.
We start from a K-way classifier that is pre-trained on inlier data $D_{\text{in}}$ as a part of an off-the-shelf model. We append the K+1-th class and fine-tune the classifier $f_{\theta}$ on the mixed-content dataset $\mathcal{D}$ by optimizing the standard cross-entropy loss:
$L_{\text{cls}}(\theta)=\mathbb{E}_{\mathbf{x},{y}\in\mathcal{D}}[-\ln P(y|\mathbf{x})]\,.$ (5)
Every training minibatch contains an equal number of samples for all K+1 classes. Thus, we expose our model to significantly less negative data than alternative approaches [Zhang et al.(2023a)Zhang, Inkawhich, Linderman, Chen, and Li, Yu and Aizawa(2019), Yang et al.(2021)Yang, Wang, Feng, Yan, Zheng, Zhang, and Liu, Hendrycks and Gimpel(2017)] that utilize 2×2$\times$2 × more negatives than inliers. Even though the outlier exposure of our model is weak, it will still be biased towards detection of the test outliers that are similar to the training negatives. To circumvent this issue, we replace the auxiliary negative dataset with synthetic negatives produced by a generative model [Lee et al.(2018)Lee, Lee, Lee, and Shin, Grcić et al.(2021)Grcić, Bevandić, and Šegvić].
Synthetic negatives. We generate model-specific synthetic negatives by a normalizing flow $f_{\psi}$ [Dinh et al.(2017)Dinh, Sohl-Dickstein, and Bengio, Kingma and Dhariwal(2018)] that we train alongside our discriminative classifier $f_{\theta}$. We train the flow to generate samples that resemble inliers, but at the same time give raise to uncertain predictions across the inlier classes according to the following loss [Lee et al.(2018)Lee, Lee, Lee, and Shin, Grcic et al.(2024)Grcic, Bevandic, Kalafatic, and Segvic]:
$L_{\text{flow}}(\psi,\theta)=L_{\text{mle}}(\psi)+\beta\cdot L_{\text{jsd}}(\psi,\theta).$ (6)
The first loss term $L_{\mathrm{mle}}(\psi)=\textbf{E}_{\mathbf{x^{+}}\in D_{\text{in}}}[-\ln p_{\psi}(\mathbf{x})]$ maximizes the likelihood of inlier samples, while the second loss term $L_{\text{jsd}}(\psi,\theta)=\text{JSD}(U,f_{\theta}(\mathbf{x}))$ minimizes the JS divergence between uniform distribution and the predictions over the K inlier classes. The two competing objectives settle down when the generative model produces samples at the border of the inlier manifold [Lee et al.(2018)Lee, Lee, Lee, and Shin]. The above loss has to jointly optimize $f_{\theta}$ and $f_{\psi}$ so that the discriminative model gets aware of the negative class and thus preclude feature collapse [Lucas et al.(2019)Lucas, Shmelkov, Alahari, Schmid, and Verbeek]. Of course, the training also has to ensure that $f_{\theta}$ satisfies the primary discriminative loss.
However, a naive application of this recipe in the K+1-way context would not work as intended. Infact, the generative loss (6) and the discriminative loss (5) would simply be satisfied by confident classification of all generated samples into the negative class. This trivial solution would embed the generated samples into a compact region of the feature space, and make them useless as proxies for outlier detection. Thus, we propose a two-step optimization procedure. The first step jointly trains $f_{\psi}$, $f_{\theta}$, and the K inlier logits according to (6). The second step freezes the flow and fine-tunes the loss (5) in order to ensure recognition of generated negatives as the K+1-th class. Different from previous approaches [Lee et al.(2018)Lee, Lee, Lee, and Shin, Grcic et al.(2024)Grcic, Bevandic, Kalafatic, and Segvic], the loss (6) serves only to promote generation of useful synthetic negatives, while we ensure the outlier recognition quality by subsequent fine-tuning according to the loss (5).
Figure 3.a illustrates the learned latent representations as obtained by the two-step training procedure. Our synthetic negatives are sprinkled around the inlier manifold. Joint optimisation of the loss (6) and the K+1-way cross-entropy loss (5) collapses synthetic negatives to a single mode, as shown in Figure 3.b. For reference, we also show the latent representations of the model trained with real negatives in Figure 3.c.
Figure 3: Left: Visualization of real and synthetic negatives from an OpenOOD CIFAR-10 experiment [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al.]. Right: t-SNE plots of the corresponding feature representations. Our two-set training strategy yields synthetic negatives near the inlier manifold (a), while the naive approach collapses synthetic negatives to a single mode (b). Relative location of real negatives indicate that they cover similar modes of test outliers as our synthetic samples (c).
Figure 4 shows the scaling of our training approach to dense recognition. We paste negative training content atop the regular image to produce a mixed-content image. The mixed-content image is fed to K+1-way classifier that optimizes the cross-entropy objective, similar to the image-wide setup. In the case of synthetic negatives, we jointly optimize the dense classifier and the normalizing flow.
Figure 4: Fine-tuning of dense classifier equipped with UNO. We paste negative training data (either real or synthetic) atop regular inlier images. The resulting mixed-content image is fed to the dense classifier that optimizes cross-entropy loss over K+1 classes.
5 Experiments
Experimental setup. We evaluate the performance of UNO on standard image-wide and pixel-level outlier detection benchmarks. In the pixel-level setup, we consider Fishyscapes [Blum et al.(2021)Blum, Sarlin, Nieto, Siegwart, and Cadena] and SMIYC [Chan et al.(2021a)Chan, Lis, Uhlemeyer, Blum, Honari, Siegwart, Fua, Salzmann, and Rottmann], the two prominent benchmarks for semantic anomaly segmentation in road driving scenes. In the case of image-wide setup, we consider the CIFAR-10 and ImageNet-200 setups from the OpenOOD benchmark [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al.]. We report the standard evaluation metrics AUROC and FPR95 in image-wide, and AP and FPR95 in dense prediction experiments. Our pixel-level experiments start by training the M2F-SwinL [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar] on Cityscapes [Cordts et al.(2016)Cordts, Omran, Ramos, Rehfeld, Enzweiler, Benenson, Franke, Roth, and Schiele] and Vistas [Neuhold et al.(2017)Neuhold, Ollmann, Rota Bulo, and Kontschieder] with the Cityscapes taxonomy. We extend the mask-wide classifier to K+2 classes and fine-tune on mixed-content images. We compose mixed-content images by pasting negative content (ADE20k [Zhou et al.(2019)Zhou, Zhao, Puig, Xiao, Fidler, Barriuso, and Torralba] instances or flow samples) into inlier images. More implementation details are in the Appendix.
Segmentation of road scenes. Table 1 compares UNO with the related work on Fishyscapes and SMIYC. We observe that mask-level recognition methods outperform earlier works [Jung et al.(2021)Jung, Lee, Gwak, Choi, and Choo, Grcic et al.(2022)Grcic, Bevandic, and Segvic, Vojir et al.(2021)Vojir, Šipka, Aljundi, Chumerin, Reino, and Matas]. When training with real negatives, our method outperforms all previous work on Fishyscapes and SMIYC AnomalyTrack while performing within variance of the best method on SMIYC ObstacleTrack. When training without real negatives, our method outperforms all previous work on SMIYC and Fishyscapes Static by a large margin while achieving second best AP on Fishyscapes L&F which appears to be a very challenging test benchmark for mask-level recognition methods because it includes small anomalies.
Fishyscapes SMIYC Cityscapes
Aux. Lost&Found Static AnomalyTrack ObstacleTrack val
Method data AP $\text{FPR}_{95}$ AP $\text{FPR}_{95}$ AP $\text{FPR}_{95}$ AP $\text{FPR}_{95}$ mIoU
Maximum Entropy [Chan et al.(2021b)Chan, Rottmann, and Gottschalk] 15.0 85.1 0.8 77.9 - - - - 9.7
Image Resynthesis [Lis et al.(2019)Lis, Nakka, Fua, and Salzmann] 5.7 48.1 29.6 27.1 52.3 25.9 37.7 4.7 81.4
JSRNet [Vojir et al.(2021)Vojir, Šipka, Aljundi, Chumerin, Reino, and Matas] - - - - 33.6 43.9 28.1 28.9 -
Max softmax [Hendrycks and Gimpel(2016)] 1.8 44.9 12.9 39.8 28.0 72.1 15.7 16.6 80.3
SML [Jung et al.(2021)Jung, Lee, Gwak, Choi, and Choo] 31.7 21.9 52.1 20.5 - - - - -
Embedding Density [Blum et al.(2021)Blum, Sarlin, Nieto, Siegwart, and Cadena] 37.5 70.8 0.8 46.4 4.3 47.2 62.1 17.4 80.3
NFlowJS [Grcić et al.(2021)Grcić, Bevandić, and Šegvić] 39.4 9.0 52.1 15.4 - - - - 77.4
SynDHybrid [Grcić and Šegvić(2023)] 51.8 11.5 54.7 15.5 - - - - 79.9
cDNP [Galesso et al.(2022)Galesso, Argus, and Brox] 62.2 8.9 - - 88.9 11.4 72.70 1.40 -
$\text{EAM}^{\dagger}$ [Grcić et al.(2023)Grcić, Šarić, and Šegvić] 9.4 41.5 76.0 10.1 76.3 93.9 66.9 17.9 83.5
$\text{RbA}^{\dagger}$ [Nayal et al.(2023)Nayal, Yavuz, Henriques, and Güney] - - - - 86.1 15.9 87.8 3.3 -
$\text{Maskomaly}^{\dagger}$ [Ackermann et al.(2023)Ackermann, Sakaridis, and Yu] - - - - 93.4 6.9 - - -
$\text{UNO}^{\dagger}$ (ours) 56.4 55.1 91.1 1.5 96.1 2.3 89.0 0.6 83.5
SynBoost [Di Biase et al.(2021)Di Biase, Blum, Siegwart, and Cadena] 43.2 15.8 72.6 18.8 56.4 61.9 71.3 3.2 81.4
OOD Head [Bevandić et al.(2022)Bevandić, Krešo, Oršić, and Šegvić] 30.9 22.2 84.0 10.3 - - - - 77.3
Void Classifier [Blum et al.(2021)Blum, Sarlin, Nieto, Siegwart, and Cadena] 10.3 22.1 45.0 19.4 36.6 63.5 10.4 41.5 70.4
Dirichlet prior [Malinin and Gales(2018)] 34.3 47.4 84.6 30.0 - - - - 70.5
DenseHybrid [Grcic et al.(2022)Grcic, Bevandic, and Segvic] 43.9 6.2 72.3 5.5 78.0 9.8 78.7 2.1 81.0
PEBAL [Tian et al.(2022)Tian, Liu, Pang, Liu, Chen, and Carneiro] 44.2 7.6 92.4 1.7 49.1 40.8 5.0 12.7 -
cDNP [Galesso et al.(2022)Galesso, Argus, and Brox] 69.8 7.5 - - 88.9 11.4 72.70 1.40 -
RPL [Liu et al.(2023)Liu, Ding, Tian, Pang, Belagiannis, Reid, and Carneiro] 53.9 2.3 95.9 0.5 83.5 11.7 85.9 0.6 -
[…]
$\text{RbA}^{\dagger}$ [Nayal et al.(2023)Nayal, Yavuz, Henriques, and Güney] - - - - 90.9 11.6 91.8 0.5 -
$\text{EAM}^{\dagger}$ [Grcić et al.(2023)Grcić, Šarić, and Šegvić] 63.5 39.2 93.6 1.2 93.8 4.1 92.9 0.5 83.5
$\text{UNO}^{\dagger}$ (ours) 74.8 2.7 95.8 0.3 96.3 2.0 93.2 0.2 83.7
Table 1: Experimental evaluation on the Fishyscapes and SMIYC benchmarks. Methods that leverage mask-level recognition are marked with $\dagger$. Missing results are marked with -.
Table 2 compares our method with other mask-level methods on validation subsets of Fishyscapes and RoadAnomaly [Lis et al.(2019)Lis, Nakka, Fua, and Salzmann]. Our method outperforms all related approaches on all datasets while only having 0.5 pp worse $\text{FPR}_{95}$ than RbA on RoadAnomaly.
FS L&F FS Static RoadAnomaly FS L&F FS Static RoadAnomaly
Method AP $\text{FPR}_{95}$ AP $\text{FPR}_{95}$ AP $\text{FPR}_{95}$ AP $\text{FPR}_{95}$ AP $\text{FPR}_{95}$ AP $\text{FPR}_{95}$
RbA 70.8 6.3 - - 85.4 6.9 61.0 10.6 - - 78.5 11.8
Maskomaly 69.4 9.4 90.5 2.0 79.7 13.5 - - 68.8 15.0 80.8 12.0
EAM 81.5 4.2 96.0 0.3 69.4 7.7 52.0 20.5 87.3 2.1 66.7 13.4
UNO (ours) 81.8 1.3 98.0 0.04 88.5 7.4 74.5 6.9 96.9 0.1 82.4 9.2
Table 2: Validation of mask-level approaches on Fishyscapes val and RoadAnomaly with (left) and without (right) training with real negative data. Missing results are marked with -.
Image classification. Table 3 shows the image-wide performance of UNO on the OpenOOD benchmark [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al.]. We compare UNO with training methods with and without the use of real negative data. When training with real negatives UNO consistently outperforms all baselines on CIFAR-10 and near OOD Imagenet-200 while attaining the best AUROC and the second-best FPR on the far OOD Imagenet-200. When training without real negative data UNO accomplishes competitive performance on both benchmarks. Note that UNO is indeed a training regularisation method, but with data as the regularisation source contrary to specially designed training algorithms. Synthetic UNO outperforms even methods trained with real data, specifically, MixOE, MCD and UDG on CIFAR-10 and UDG on ImageNet-200. It is interesting that UNO trained with synthetic negatives outperforms UNO trained with real negatives by a significant margin on the far OOD ImageNet-200. Finally, UNO preserves the inlier classification performance.
Method CIFAR-10 ImageNet-200
Near-OOD Far-OOD Near-OOD Far-OOD
AUC FPR AUC FPR Acc. AUC FPR AUC FPR Acc.
ConfBranch [DeVries and Taylor(2018)] 89.8 31.3 92.9 94.9 94.9 79.1 61.4 90.4 34.8 85.9
RotPred [Hendrycks et al.(2019b)Hendrycks, Mazeika, Kadavath, and Song] 92.7 28.1 96.6 12.2 95.4 81.6 60.4 92.6 26.2 86.4
G-ODIN [Hsu et al.(2020)Hsu, Shen, Jin, and Kira] 89.1 45.5 95.5 21.5 94.7 77.3 69.9 92.3 30.2 84.6
CSI [Tack et al.(2020)Tack, Mo, Jeong, and Shin] 89.5 33.7 92.0 26.4 91.2 - - - - -
ARPL [Chen et al.(2022)Chen, Peng, Wang, and Tian] 87.4 40.3 89.3 32.4 93.7 82.0 55.7 89.2 36.5 84.0
MOS [Huang and Li(2021)] 71.5 78.7 76.4 62.9 94.8 69.8 71.6 80.5 51.6 85.6
VOS [Du et al.(2022)Du, Wang, Cai, and Li] 87.7 57.0 90.8 40.4 94.3 82.5 59.9 91.0 34.0 86.2
LogitNorm [Wei et al.(2022)Wei, Xie, Cheng, Feng, An, and Li] 92.3 29.3 96.7 13.8 94.3 82.7 56.5 93.0 26.1 86.0
CIDER [Ming et al.(2023)Ming, Sun, Dia, and Li] 90.7 32.1 94.7 20.7 - 80.6 60.1 90.7 30.2 -
NPOS [Tao et al.(2023)Tao, Du, Zhu, and Li] 89.8 32.6 94.1 20.6 - 79.4 62.1 94.5 21.8 -
UNO (ours) 91.3 31.8 92.6 20.5 95.2 81.2 61.1 92.5 32.3 86.3
MixOE [Zhang et al.(2023a)Zhang, Inkawhich, Linderman, Chen, and Li] 88.7 51.5 91.9 33.8 94.6 82.6 58.0 88.3 40.9 85.7
MCD [Yu and Aizawa(2019)] 91.0 30.2 91.0 32.0 95.0 83.6 54.7 88.9 29.9 86.1
UDG [Yang et al.(2021)Yang, Wang, Feng, Yan, Zheng, Zhang, and Liu] 89.9 35.3 92.4 20.4 92.4 74.3 68.9 82.1 62.0 68.1
OE [Hendrycks and Gimpel(2017)] 94.8 19.8 96.0 13.1 94.6 84.8 52.3 89.0 34.2 85.8
UNO (ours) 94.9 19.3 97.6 9.4 94.9 85.1 51.7 89.6 36.8 86.4
Table 3: OOD detection performance on OpenOOD CIFAR-10 and ImageNet-200. We consider training-based methods with (bottom) and without (top) real negative data. All results were averaged over 3 runs. Full results are available in the Appendix.
Validating components of UNO score. Table 4 validates the contribution of the UNO components. In the case of real negatives, both components attain competitive results while UNO significantly outperforms both of them. In the case of synthetic negatives, uncertainty outperforms negative objectness by a wide margin. This does not come as a surprise since the negative objectness is inferred from synthetic training negatives. Still, UNO greatly benefits from the ensemble of the two scores.
Score FS L&F FS Static RoadAnomaly FS L&F FS Static RoadAnomaly
AP $\text{FPR}_{95}$ AP $\text{FPR}_{95}$ AP $\text{FPR}_{95}$ AP $\text{FPR}_{95}$ AP $\text{FPR}_{95}$ AP $\text{FPR}_{95}$
UNO 81.8 1.3 98.0 0.04 88.5 7.4 74.5 6.9 96.9 0.1 82.4 9.2
Unc 74.1 4.5 72.1 1.5 66.2 8.0 71.9 8.0 95.7 0.5 70.4 9.4
NO 69.0 1.6 92.6 0.14 80.4 19.8 26.6 91.1 73.9 61.1 54.2 72.3
Table 4: Validation of UNO components with (left) and without (right) real negative data.
Figure 5 visualises UNO performance on FS L&F. Columns show continuous OOD scores and OOD detections after thresholding at 95% TPR. The two UNO components have different failure modes. For example, $\textbf{s}_{\text{Unc}}$ produces false positives at the borders of inlier classes while $\textbf{s}_{\text{NO}}$ detects some inlier objects as outliers. Still, these failure modes cancel out in the compound UNO score, as designated with green rectangles.
Figure 5: Qualitative experiments on Fishyscapes L&F val. Top row shows the input image and the three outlier scores. Bottom row shows anomaly detection maps after thresholding at TPR=95%. UNO significantly reduces false-positive responses.
6 Conclusion
Visual recognition models behave unpredictably in presence of outliers. A common strategy to detect outliers is to fine-tune the closed-set classifier in order to increase uncertainty in negative data. We take a different approach and disentangle the prediction uncertainty from the recognition of outliers. This allows us to cast outlier detection as an ensemble of in-distribution uncertainty and the posterior of the negative class, which we term negative objectness. The resulting UNO score manifests as a lightweight plug-in module that can extend an arbitrary classifier. The resulting model is then fine-tuned with real or synthetic negative data that imitate test outliers. UNO can be applied to per-pixel outlier detection by extending recent mask-level recognition approaches. There we first apply UNO to recover mask-wide outlier scores, and then propagate them to pixels according to mask assignments. UNO attains strong performance on recent image-wide and pixel-level outlier detection benchmarks with and without real negative data. Prominent future developments include improving the quality of synthetic negative data and extending our ensemble with generative predictions. Limitations of UNO are discussed in the Appendix.
Acknowledgements
This work has been co-funded by European Union - NextGenerationEU and Croatian Science Foundation grant IP-2020-02-5851 ADEPT.

### v4

@@
 Although it cannot represent the entire variety of the visual world, the negative data can still help by signaling that not all data should be confidently recognized.
-If the variety of the negative data exceeds the variety of the inliers, then there is a reasonable hope that the test outliers will be detected.
+If the variety of the negative data greatly exceeds the variety of the inliers, then there is a reasonable hope that the test outliers will be detected.
 Unfortunately, this entails undesired bias towards test outliers that appear similar to the negative training data.
-This concern can be addressed either by relying on synthetic negatives [Lee et al.(2018)Lee, Lee, Lee, and Shin, Neal et al.(2018)Neal, Olson, Fern, Wong, and Li], or through separate ranking with respect to approaches that do not train on real negative data [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al., Blum et al.(2021)Blum, Sarlin, Nieto, Siegwart, and Cadena, Chan et al.(2021a)Chan, Lis, Uhlemeyer, Blum, Honari, Siegwart, Fua, Salzmann, and Rottmann].
+This concern can be addressed either by relying on synthetic negatives [Lee et al.(2018a)Lee, Lee, Lee, and Shin, Neal et al.(2018)Neal, Olson, Fern, Wong, and Li, Grcić and Šegvić(2024b)], or through separate ranking with respect to approaches that do not train on real negative data [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al., Blum et al.(2021)Blum, Sarlin, Nieto, Siegwart, and Cadena, Chan et al.(2021a)Chan, Lis, Uhlemeyer, Blum, Honari, Siegwart, Fua, Salzmann, and Rottmann].
 Figure 1: We propose UNO, a plug-in module that enables outlier-aware inference atop a desired feature extractor.
 UNO decouples in-distribution uncertainty from outlier recognition, and boosts the outlier detection performance by ensembling the two components.
-Many existing methods improve open-set performance of existing closed-set models by encouraging low-confidence predictions in negative training data [Hendrycks et al.(2022)Hendrycks, Basart, Mazeika, Zou, Kwon, Mostajabi, Steinhardt, and Song, Lee et al.(2018)Lee, Lee, Lee, and Shin].
+Many existing methods improve open-set performance of existing closed-set models by encouraging low-confidence predictions in negative training data [Hendrycks et al.(2022a)Hendrycks, Basart, Mazeika, Zou, Kwon, Mostajabi, Steinhardt, and Song, Lee et al.(2018a)Lee, Lee, Lee, and Shin].
 However, this approach conflates prediction uncertainty with outlier recognition.
-In this paper, we disentangle the two factors by reconsidering the K+1-way classifier where the additional class represents the training negative data.
+In this paper, we disentangle the two factors by reconsidering the K+1-way classifier where the additional class represents the negative data.
 This allows us to formulate the outlier score as the posterior of the K+1-th class, which we term negative objectness.
-Moreover, the outlier score can also be formulated as prediction uncertainty over the K known classes [Hendrycks et al.(2022)Hendrycks, Basart, Mazeika, Zou, Kwon, Mostajabi, Steinhardt, and Song].
+Moreover, the outlier score can also be formulated as prediction uncertainty over the K known classes [Hendrycks et al.(2022a)Hendrycks, Basart, Mazeika, Zou, Kwon, Mostajabi, Steinhardt, and Song].
 We find that these two formulations exhibit complementary behaviour due to being sensitive to different test outliers.
 Consequently, we propose a novel outlier score that we term UNO: ensemble of Uncertainty over K inlier classes [Hendrycks and Gimpel(2017)] and Negative Objectness.
-Our UNO score is a strikingly good fit in the pixel-level context as an extension of the direct set-prediction approach [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar, Yu et al.(2022)Yu, Wang, Qiao, Collins, Zhu, Adam, Yuille, and Chen, Li et al.(2023)Li, Zhang, xu, Liu, Zhang, Ni, and Shum].
-We express the negative objectness as the posterior of the negative class in the frame of a K+2-way mask-level classifier.
+Our UNO score is a strikingly good fit in the pixel-level context as an extension of some direct set-prediction approach [Carion et al.(2020)Carion, Massa, Synnaeve, Usunier, Kirillov, and Zagoruyko].
+We express the negative objectness as the posterior of the negative class in the frame of a K+2-way mask-level classifier [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar, Yu et al.(2022)Yu, Wang, Qiao, Collins, Zhu, Adam, Yuille, and Chen, Li et al.(2023)Li, Zhang, xu, Liu, Zhang, Ni, and Shum].
 A closer look suggests that our score performs well due to different inductive bias of the two components as indicated by a remarkably weak correlation.
@@
 2 Related work
-Image-wide classification.
+Image-wide setup.
 Image-wide classification models can be complemented with outlier detection either through a modified training procedure or in a post-hoc manner with frozen parameters [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al.].
@@
 TempScale [Guo et al.(2017)Guo, Pleiss, Sun, and Weinberger] calibrates softmax probabilities with temperature scaling while ODIN [Liang et al.(2017)Liang, Li, and Srikant] pre-processes the input with anti-adversarial perturbations.
-Recent post-hoc methods simplify layer activations [Djurisic et al.(2022)Djurisic, Bozanic, Ashok, and Liu] or consider distances from class prototypes [Zhang et al.(2022)Zhang, Fu, Chen, Du, Li, Wang, Han, Zhang, et al.].
+Recent post-hoc methods simplify layer activations [Djurisic et al.(2022)Djurisic, Bozanic, Ashok, and Liu] or fit generative models to pre-logit features [Zhang et al.(2022)Zhang, Fu, Chen, Du, Li, Wang, Han, Zhang, et al., Yu et al.(2022)Yu, Wang, Qiao, Collins, Zhu, Adam, Yuille, and Chen, Taghikhah et al.(2024)Taghikhah, Kumar, Segvic, Eslami, and Gumhold].
 In contrast, training methods involve various regularizations that either consider only inlier data or utilize real negative data as well.
-The former rely on modeling confidence [DeVries and Taylor(2018)], self-supervised training [Hendrycks et al.(2019b)Hendrycks, Mazeika, Kadavath, and Song, Tack et al.(2020)Tack, Mo, Jeong, and Shin], per-sample temperature scaling [Hsu et al.(2020)Hsu, Shen, Jin, and Kira], outlier synthesis in feature space [Du et al.(2022)Du, Wang, Cai, and Li, Tao et al.(2023)Tao, Du, Zhu, and Li] or mitigating overconfidence [Wei et al.(2022)Wei, Xie, Cheng, Feng, An, and Li].
+The former rely on modeling confidence [DeVries and Taylor(2018)], self-supervised training [Hendrycks et al.(2019b)Hendrycks, Mazeika, Kadavath, and Song, Tack et al.(2020)Tack, Mo, Jeong, and Shin], per-sample temperature scaling [Hsu et al.(2020)Hsu, Shen, Jin, and Kira], outlier synthesis in feature space [Du et al.(2022)Du, Wang, Cai, and Li, Tao et al.(2023)Tao, Du, Zhu, and Li, Kumar et al.(2023)Kumar, Segvic, Eslami, and Gumhold] or mitigating overconfidence [Wei et al.(2022)Wei, Xie, Cheng, Feng, An, and Li].
 Recent methods guide the learning process with respect to nearest neighbours [Park et al.(2023)Park, Jung, and Teoh] or leverage distances in feature space either by combining different distance functions [Noh et al.(2023)Noh, Jeong, and Lee] or distances in different feature spaces [Chen et al.(2023)Chen, Lin, Xu, and Vela].
-Early work in training with negative data [Hendrycks et al.(2018)Hendrycks, Mazeika, and Dietterich, Dhamija et al.(2018)Dhamija, Günther, and Boult] encourages classifiers to produce high entropy in inlier samples.
+Early work in training with negative data [Hendrycks et al.(2018)Hendrycks, Mazeika, and Dietterich, Dhamija et al.(2018)Dhamija, Günther, and Boult] encourages high entropy in inlier samples.
 MCD [Yu and Aizawa(2019)] promotes discrepancy between prediction on OOD samples of two classification heads.
@@
 MixOE [Zhang et al.(2023a)Zhang, Inkawhich, Linderman, Chen, and Li] interpolates between inliers and negatives to improve regularization.
-Unlike related training based methods UNO does not modify the standard classifier training objective, but merely requires a slight architectural change to extend the classifier with a logit for the negative class.
-Additionally, UNO requires significantly less negative data than related methods.
+Unlike related training-based methods, UNO does not modify the standard classifier training objective, but merely requires a slight architectural change to extend the classifier with a logit for the negative class.
+Moreover, empirical insights suggest that UNO requires significantly less negative data than related methods.
 Dense prediction.
@@
 Early approaches estimate the prediction uncertainty with maximum softmax probability [Hendrycks and Gimpel(2017)], ensembling [Lakshminarayanan et al.(2017)Lakshminarayanan, Pritzel, and Blundell] or Bayesian uncertainty [Mukhoti and Gal(2018)].
-Subsequent work suggests using energy scores [Liu et al.(2020)Liu, Wang, Owens, and Li] or standardized max-logit [Jung et al.(2021)Jung, Lee, Gwak, Choi, and Choo].
-Training with negative data can be incorporated by pasting negative content from broad negative datasets (usually ADE20K [Zhou et al.(2017)Zhou, Zhao, Puig, Fidler, Barriuso, and Torralba] or COCO [Lin et al.(2014a)Lin, Maire, Belongie, Hays, Perona, Ramanan, Dollár, and Zitnick]) atop inlier images [Biase et al.(2021)Biase, Blum, Siegwart, and Cadena, Chan et al.(2021b)Chan, Rottmann, and Gottschalk, Grcic et al.(2022)Grcic, Bevandic, and Segvic, Hendrycks et al.(2018)Hendrycks, Mazeika, and Dietterich, Tian et al.(2022)Tian, Liu, Pang, Liu, Chen, and Carneiro, Bevandić et al.(2022)Bevandić, Krešo, Oršić, and Šegvić].
-However, real negative data can be replaced with synthetic negative patches obtained by sampling a jointly trained generative model [Lee et al.(2018)Lee, Lee, Lee, and Shin, Neal et al.(2018)Neal, Olson, Fern, Wong, and Li, Zhao et al.(2021)Zhao, Cao, and Lin].
+Subsequent work suggests energy scores [Liu et al.(2020)Liu, Wang, Owens, and Li] and the standardized max-logit score [Jung et al.(2021)Jung, Lee, Gwak, Choi, and Choo].
+Training with negative data can be incorporated by pasting negative content from broad negative datasets (ImageNet [Bevandić et al.(2019)Bevandić, Krešo, Oršić, and Šegvić], ADE20K [Zhou et al.(2017b)Zhou, Zhao, Puig, Fidler, Barriuso, and Torralba] or COCO [Lin et al.(2014a)Lin, Maire, Belongie, Hays, Perona, Ramanan, Dollár, and Zitnick]) atop inlier images [Biase et al.(2021)Biase, Blum, Siegwart, and Cadena, Chan et al.(2021b)Chan, Rottmann, and Gottschalk, Grcic et al.(2022)Grcic, Bevandic, and Segvic, Hendrycks et al.(2018)Hendrycks, Mazeika, and Dietterich, Tian et al.(2022)Tian, Liu, Pang, Liu, Chen, and Carneiro, Bevandić et al.(2022)Bevandić, Krešo, Oršić, and Šegvić].
+However, real negative data can be replaced with synthetic negative patches obtained by sampling a jointly trained generative model [Lee et al.(2018a)Lee, Lee, Lee, and Shin, Neal et al.(2018)Neal, Olson, Fern, Wong, and Li, Zhao et al.(2021)Zhao, Cao, and Lin].
 Outlier detection performance can be further improved by providing more capacity [Vaze et al.(2021)Vaze, Han, Vedaldi, and Zisserman].
-Another branch of outlier detection methods considers generative models [Blum et al.(2021)Blum, Sarlin, Nieto, Siegwart, and Cadena, Grcic et al.(2022)Grcic, Bevandic, and Segvic, Liang et al.(2022)Liang, Wang, Miao, and Yang].
-Mask-level recognition Recent panoptic architectures [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar, Yu et al.(2022)Yu, Wang, Qiao, Collins, Zhu, Adam, Yuille, and Chen, Li et al.(2023)Li, Zhang, xu, Liu, Zhang, Ni, and Shum] decompose scene understanding into class-agnostic segmentation and region-wide recognition.
+Another branch of methods considers generative models [Blum et al.(2021)Blum, Sarlin, Nieto, Siegwart, and Cadena, Grcic et al.(2022)Grcic, Bevandic, and Segvic, Liang et al.(2022)Liang, Wang, Miao, and Yang].
+Mask-level recognition.
+Recent panoptic architectures [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar, Yu et al.(2022)Yu, Wang, Qiao, Collins, Zhu, Adam, Yuille, and Chen, Li et al.(2023)Li, Zhang, xu, Liu, Zhang, Ni, and Shum] decompose scene understanding into class-agnostic segmentation and region-wide recognition.
 They frame the detection of semantic masks as direct set-prediction where each mask is classified into K inlier classes and one no-object class.
@@
 Let $h_{\theta_{1}}:\mathcal{X}\rightarrow\mathbb{R}^{d}$ be a feature extractor (e.g.
-ResNet [He et al.(2016)He, Zhang, Ren, and Sun] or ViT [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby]) onto which we attach our K+1-way classifier $g_{\theta_{2}}:\mathbb{R}^{d}\rightarrow\Delta^{K}$, where $\Delta^{K}$ is a K-dimensional probabilistic simplex.
+ResNet [He et al.(2016)He, Zhang, Ren, and Sun] or ViT [Dosovitskiy et al.(2021)Dosovitskiy, Beyer, Kolesnikov, Weissenborn, Zhai, Unterthiner, Dehghani, Minderer, Heigold, Gelly, Uszkoreit, and Houlsby]) onto which we attach our K+1-way classifier $g_{\theta_{2}}:\mathbb{R}^{d}\rightarrow\Delta^{K}$, where $\mathcal{X}$ is the input space and $\Delta^{K}$ is a K-dimensional probabilistic simplex.
 Given an input $\mathbf{x}\in\mathcal{X}$, we compute a latent representation $\mathbf{z}=h_{\theta_{1}}(\mathbf{x})$ and map it onto the simplex point $\mathbf{p}=\text{softmax}(g_{\theta_{2}}(\mathbf{z}))$.
@@
 Thus we have $g_{\theta_{2}}(\textbf{x})=\textbf{W}\cdot\textbf{x}$, where the rows of W correspond to class vectors $\textbf{w}_{j}$.
-Negative objectness. Given feature representation $\mathbf{z}$ and a linear classifier $g$, we formulate our negative objectness score ($\textbf{s}_{\text{NO}}$) as posterior of the K+1-th class:
+Negative objectness.
+Given feature representation $\mathbf{z}$ and a linear classifier $g$, we formulate the negative objectness score ($\textbf{s}_{\text{NO}}$) as posterior of the K+1-th class:
 $\mathbf{s}_{\text{NO}}(\mathbf{z}):=P(Y=\text{K+1}|\mathbf{z})=\frac{\text{exp}(\mathbf{w}_{K+1}^{T}\mathbf{z})}{\sum_{j=1}^{K+1}\text{exp}(\mathbf{w}_{j}^{T}\mathbf{z})}\;.$ (1)
@@
 Prediction uncertainty.
-Negative objectness is not suitable for detecting test outliers that are semantically dissimilar from the training data.
+Negative objectness is not suitable for detecting test outliers that are semantically dissimilar from the negative training data.
 However, many of these outliers yield uncertain predictions across inlier classes of our K+1-classifier.
@@
 The uncertainty score $\textbf{s}_{\text{Unc}}$ detects different outliers than negative objectness $\textbf{s}_{\text{NO}}$ despite attending to the same shared features $\mathbf{z}$.
-UNO score. We integrate the prediction uncertainty (2) and negative objectness (1) into the UNO score as the sum of the two terms:
+UNO score.
+We integrate the prediction uncertainty (2) and negative objectness (1) into the UNO score as the sum of the two terms:
 $\mathbf{s}_{\text{UNO}}(\mathbf{z})=\mathbf{s}_{\text{Unc}}(\mathbf{z})+\mathbf{s}_{\text{NO}}(\mathbf{z}).$ (3)
 UNO has an interesting geometrical interpretation in the pre-logit space $\mathbb{R}^{d}$.
-We note that class vectors tend to be mutually orthogonal: $\mathbf{w}_{i}^{T}\mathbf{w}_{j}=0,\,\,\forall i,j\in\{1,2,...,K+1\}$.
-This arrangement is enforced by the standard supervised loss ($-\log\text{softmax}$) whenever there is enough data, as well as enough capacity and dimensionality ($d>K+1$) in the feature extractor $h_{\theta_{1}}$.
+Class vectors tend to be mutually orthogonal: $\mathbf{w}_{i}^{T}\mathbf{w}_{j}=0,\,\,\forall i,j\in\{1,2,...,K+1\}$ [Wang et al.(2022)Wang, Li, Feng, and Zhang].
+This arrangement is enforced by the standard supervised loss ($-\log\text{softmax}$) whenever there is enough data, and enough capacity and dimensionality ($d>K+1$) in the feature extractor $h_{\theta_{1}}$.
 Thus, a small angle between the latent representation z and the negative vector $\textbf{w}_{K+1}$ leads to high outlier objectness $\textbf{s}_{\text{NO}}$.
-Moreover, we note that the uncertainty score $\textbf{s}_{\text{Unc}}$ is negatively correlated with the feature norm $\lvert\lvert\mathbf{z}\rvert\rvert$.
+On the other hand, the uncertainty score $\textbf{s}_{\text{Unc}}$ is negatively correlated with the feature norm $\lvert\lvert\mathbf{z}\rvert\rvert$.
 If we fix the direction of z, its norm can be viewed as reciprocal softmax temperature.
 As usual, large temperatures lead to uncertain predictions [Hinton et al.(2015)Hinton, Vinyals, and Dean].
-The above insights show that the two components of our score will be at least partially decorrelated due to being sensitive to different outliers.
+The above insights show that the two components of our score will be at least partially decorrelated due to being sensitive to different outliers: $\textbf{s}_{\text{NO}}$ detects large feature norms and good alignment with $\mathbf{w}_{K+1}$, while $\textbf{s}_{\text{Unc}}$ detects small feature norms and poor alignment with inlier vectors.
 Thus, the UNO score is likely to achieve lower error than any of the two components alone.
@@
 However, in the case of dense prediction, many contemporary architectures rely on mask-level recognition [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar].
-Consequently, we build our per-pixel score by merging multiple mask-level UNO scores:
+Consequently, we build our per-pixel score by averaging the mask-level UNO scores:
 $\textbf{s}_{\text{UNO}}^{\text{M2F}}=\sum_{i=1}^{N}\mathbf{m}_{i}\cdot\textbf{s}_{\text{UNO}}(\textbf{z}_{i})\;.$ (4)
-Equation 4 assumes that a dense prediction model has partitioned the input image into $N$ different masks $\mathbf{m}_{i}$ with the corresponding latent mask embeddings $\mathbf{z}_{i}$.
+Note that equation 4 assumes that the dense prediction model has partitioned the input image into $N$ different masks $\mathbf{m}_{i}$ with the corresponding latent mask embeddings $\mathbf{z}_{i}$.
 We note that the mask-level classifiers already use the K+1-th logit to indicate that the particular query has not been associated with any image region.
@@
 Every training minibatch contains an equal number of samples for all K+1 classes.
-Thus, we expose our model to significantly less negative data than alternative approaches [Zhang et al.(2023a)Zhang, Inkawhich, Linderman, Chen, and Li, Yu and Aizawa(2019), Yang et al.(2021)Yang, Wang, Feng, Yan, Zheng, Zhang, and Liu, Hendrycks and Gimpel(2017)] that utilize 2×2$\times$2 × more negatives than inliers.
+Thus, we expose our model to significantly less negative data than the alternative approaches [Zhang et al.(2023a)Zhang, Inkawhich, Linderman, Chen, and Li, Yu and Aizawa(2019), Yang et al.(2021)Yang, Wang, Feng, Yan, Zheng, Zhang, and Liu, Hendrycks and Gimpel(2017)], which utilize 2×2$\times$2 × more negatives than inliers.
 Even though the outlier exposure of our model is weak, it will still be biased towards detection of the test outliers that are similar to the training negatives.
-To circumvent this issue, we replace the auxiliary negative dataset with synthetic negatives produced by a generative model [Lee et al.(2018)Lee, Lee, Lee, and Shin, Grcić et al.(2021)Grcić, Bevandić, and Šegvić].
+This issue can be circumvented by replacing the auxiliary negative dataset with synthetic negatives produced by a generative model [Lee et al.(2018a)Lee, Lee, Lee, and Shin, Grcić et al.(2021b)Grcić, Bevandić, and Šegvić].
 Synthetic negatives.
 We generate model-specific synthetic negatives by a normalizing flow $f_{\psi}$ [Dinh et al.(2017)Dinh, Sohl-Dickstein, and Bengio, Kingma and Dhariwal(2018)] that we train alongside our discriminative classifier $f_{\theta}$.
-We train the flow to generate samples that resemble inliers, but at the same time give raise to uncertain predictions across the inlier classes according to the following loss [Lee et al.(2018)Lee, Lee, Lee, and Shin, Grcic et al.(2024)Grcic, Bevandic, Kalafatic, and Segvic]:
+We train the flow to generate samples that resemble inliers, but at the same time give raise to uncertain predictions across the inlier classes according to the following loss [Lee et al.(2018a)Lee, Lee, Lee, and Shin, Grcic et al.(2024)Grcic, Bevandic, Kalafatic, and Segvic]:
 $L_{\text{flow}}(\psi,\theta)=L_{\text{mle}}(\psi)+\beta\cdot L_{\text{jsd}}(\psi,\theta).$ (6)
-The first loss term $L_{\mathrm{mle}}(\psi)=\textbf{E}_{\mathbf{x^{+}}\in D_{\text{in}}}[-\ln p_{\psi}(\mathbf{x})]$ maximizes the likelihood of inlier samples, while the second loss term $L_{\text{jsd}}(\psi,\theta)=\text{JSD}(U,f_{\theta}(\mathbf{x}))$ minimizes the JS divergence between uniform distribution and the predictions over the K inlier classes.
-The two competing objectives settle down when the generative model produces samples at the border of the inlier manifold [Lee et al.(2018)Lee, Lee, Lee, and Shin].
-The above loss has to jointly optimize $f_{\theta}$ and $f_{\psi}$ so that the discriminative model gets aware of the negative class and thus preclude feature collapse [Lucas et al.(2019)Lucas, Shmelkov, Alahari, Schmid, and Verbeek].
+The first loss term $L_{\mathrm{mle}}(\psi)=\textbf{E}_{\mathbf{x^{+}}\in D_{\text{in}}}[-\ln p_{\psi}(\mathbf{x})]$ maximizes the likelihood of inlier samples, while the second loss term $L_{\text{jsd}}(\psi,\theta)=\text{JSD}(U,f_{\theta}(\mathbf{x}))$ minimizes the Jensen-Shannon divergence between the uniform distribution and the predictions over the K inlier classes.
+The two competing objectives settle down when the generative model produces samples at the border of the inlier manifold [Lee et al.(2018a)Lee, Lee, Lee, and Shin].
+The above loss jointly optimizes $f_{\theta}$ and $f_{\psi}$ so that the discriminative model get aware of the negative class and thus preclude feature collapse [Lucas et al.(2019)Lucas, Shmelkov, Alahari, Schmid, and Verbeek].
 Of course, the training also has to ensure that $f_{\theta}$ satisfies the primary discriminative loss.
@@
 The second step freezes the flow and fine-tunes the loss (5) in order to ensure recognition of generated negatives as the K+1-th class.
-Different from previous approaches [Lee et al.(2018)Lee, Lee, Lee, and Shin, Grcic et al.(2024)Grcic, Bevandic, Kalafatic, and Segvic], the loss (6) serves only to promote generation of useful synthetic negatives, while we ensure the outlier recognition quality by subsequent fine-tuning according to the loss (5).
+Different from previous approaches [Lee et al.(2018a)Lee, Lee, Lee, and Shin, Grcic et al.(2024)Grcic, Bevandic, Kalafatic, and Segvic], the loss (6) serves only to promote generation of useful synthetic negatives, while we ensure the outlier recognition quality by subsequent fine-tuning according to the loss (5).
 Figure 3.a illustrates the learned latent representations as obtained by the two-step training procedure.
@@
 Relative location of real negatives indicate that they cover similar modes of test outliers as our synthetic samples (c).
-Figure 4 shows the scaling of our training approach to dense recognition.
+Figure 4 shows our training approach in the dense prediction context.
 We paste negative training content atop the regular image to produce a mixed-content image.
-The mixed-content image is fed to K+1-way classifier that optimizes the cross-entropy objective, similar to the image-wide setup.
+The mixed-content image is fed to the K+2-way classifier that optimizes the cross-entropy objective, similar as in the image-wide context.
 In the case of synthetic negatives, we jointly optimize the dense classifier and the normalizing flow.
@@
 We paste negative training data (either real or synthetic) atop regular inlier images.
-The resulting mixed-content image is fed to the dense classifier that optimizes cross-entropy loss over K+1 classes.
+The resulting mixed-content image is fed to the dense classifier that optimizes cross-entropy loss over K+2 classes.
 5 Experiments
 Experimental setup.
-We evaluate the performance of UNO on standard image-wide and pixel-level outlier detection benchmarks.
-In the pixel-level setup, we consider Fishyscapes [Blum et al.(2021)Blum, Sarlin, Nieto, Siegwart, and Cadena] and SMIYC [Chan et al.(2021a)Chan, Lis, Uhlemeyer, Blum, Honari, Siegwart, Fua, Salzmann, and Rottmann], the two prominent benchmarks for semantic anomaly segmentation in road driving scenes.
-In the case of image-wide setup, we consider the CIFAR-10 and ImageNet-200 setups from the OpenOOD benchmark [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al.].
-We report the standard evaluation metrics AUROC and FPR95 in image-wide, and AP and FPR95 in dense prediction experiments.
+We evaluate the UNO performance on standard image-wide and pixel-level benchmarks.
+In the pixel-level setup, we consider Fishyscapes [Blum et al.(2021)Blum, Sarlin, Nieto, Siegwart, and Cadena] and SMIYC [Chan et al.(2021a)Chan, Lis, Uhlemeyer, Blum, Honari, Siegwart, Fua, Salzmann, and Rottmann], the two prominent benchmarks for road driving scenes.
+In the image-wide setup, we consider the CIFAR-10 and ImageNet-200 setups from the OpenOOD benchmark [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al.].
+We report the standard evaluation metrics AUROC and FPR95 in image-wide, and AP and FPR95 in pixel-level experiments.
 Our pixel-level experiments start by training the M2F-SwinL [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar] on Cityscapes [Cordts et al.(2016)Cordts, Omran, Ramos, Rehfeld, Enzweiler, Benenson, Franke, Roth, and Schiele] and Vistas [Neuhold et al.(2017)Neuhold, Ollmann, Rota Bulo, and Kontschieder] with the Cityscapes taxonomy.
@@
 More implementation details are in the Appendix.
+Our code is publicly available here [^1] https://github.com/matejgrcic/Open-set-M2F.
 Segmentation of road scenes.
@@
 We observe that mask-level recognition methods outperform earlier works [Jung et al.(2021)Jung, Lee, Gwak, Choi, and Choo, Grcic et al.(2022)Grcic, Bevandic, and Segvic, Vojir et al.(2021)Vojir, Šipka, Aljundi, Chumerin, Reino, and Matas].
-When training with real negatives, our method outperforms all previous work on Fishyscapes and SMIYC AnomalyTrack while performing within variance of the best method on SMIYC ObstacleTrack.
-When training without real negatives, our method outperforms all previous work on SMIYC and Fishyscapes Static by a large margin while achieving second best AP on Fishyscapes L&F which appears to be a very challenging test benchmark for mask-level recognition methods because it includes small anomalies.
+When training with real negatives, our method outperforms all previous work on Fishyscapes and SMIYC AnomalyTrack, while performing within variance of the best method on SMIYC ObstacleTrack.
+When training without real negatives, our method outperforms all previous work on SMIYC and Fishyscapes Static by a large margin, while achieving second best AP on Fishyscapes L&F.
+Interestingly, the inlier recognition performance is not affected by the UNO fine-tuning.
+The initial closed-set model attains 83.5% mIoU on Cityscapes.
+Appending UNO and fine-tuning the resulting open-set model with real negatives increases classification performance to 83.7%.
 Fishyscapes SMIYC Cityscapes
@@
 Embedding Density [Blum et al.(2021)Blum, Sarlin, Nieto, Siegwart, and Cadena] 37.5 70.8 0.8 46.4 4.3 47.2 62.1 17.4 80.3
-NFlowJS [Grcić et al.(2021)Grcić, Bevandić, and Šegvić] 39.4 9.0 52.1 15.4 - - - - 77.4
-SynDHybrid [Grcić and Šegvić(2023)] 51.8 11.5 54.7 15.5 - - - - 79.9
+NFlowJS [Grcić et al.(2021b)Grcić, Bevandić, and Šegvić] 39.4 9.0 52.1 15.4 - - - - 77.4
+SynDHybrid [Grcić and Šegvić(2024a)] 51.8 11.5 54.7 15.5 - - - - 79.9
 cDNP [Galesso et al.(2022)Galesso, Argus, and Brox] 62.2 8.9 - - 88.9 11.4 72.70 1.40 -
@@
 Table 1: Experimental evaluation on the Fishyscapes and SMIYC benchmarks. Methods that leverage mask-level recognition are marked with $\dagger$. Missing results are marked with -.
-Table 2 compares our method with other mask-level methods on validation subsets of Fishyscapes and RoadAnomaly [Lis et al.(2019)Lis, Nakka, Fua, and Salzmann].
-Our method outperforms all related approaches on all datasets while only having 0.5 pp worse $\text{FPR}_{95}$ than RbA on RoadAnomaly.
+Table 2 compares UNO with previous works that work with mask-level recognition on validation subsets of Fishyscapes and RoadAnomaly [Lis et al.(2019)Lis, Nakka, Fua, and Salzmann].
+UNO outperforms previous approaches across datasets and metrics with the single exception of 0.5 pp worse $\text{FPR}_{95}$ than RbA on RoadAnomaly.
 FS L&F FS Static RoadAnomaly FS L&F FS Static RoadAnomaly
@@
 We compare UNO with training methods with and without the use of real negative data.
-When training with real negatives UNO consistently outperforms all baselines on CIFAR-10 and near OOD Imagenet-200 while attaining the best AUROC and the second-best FPR on the far OOD Imagenet-200.
-When training without real negative data UNO accomplishes competitive performance on both benchmarks.
-Note that UNO is indeed a training regularisation method, but with data as the regularisation source contrary to specially designed training algorithms.
-Synthetic UNO outperforms even methods trained with real data, specifically, MixOE, MCD and UDG on CIFAR-10 and UDG on ImageNet-200.
+When training with real negatives, UNO consistently outperforms all baselines on CIFAR-10 and near OOD Imagenet-200, while attaining the best AUROC and the second-best FPR on the far OOD Imagenet-200.
+When training without real negative data, UNO accomplishes competitive performance on both benchmarks.
+Synthetic UNO outperforms even methods trained with real data, specifically, MixOE, MCD and UDG on CIFAR-10, and UDG on ImageNet-200.
 It is interesting that UNO trained with synthetic negatives outperforms UNO trained with real negatives by a significant margin on the far OOD ImageNet-200.
@@
 OE [Hendrycks and Gimpel(2017)] 94.8 19.8 96.0 13.1 94.6 84.8 52.3 89.0 34.2 85.8
-UNO (ours) 94.9 19.3 97.6 9.4 94.9 85.1 51.7 89.6 36.8 86.4
+UNO (ours) 94.9 9.3 97.6 9.4 94.9 85.1 51.7 89.6 36.8 86.4
 Table 3: OOD detection performance on OpenOOD CIFAR-10 and ImageNet-200.
@@
 In the case of real negatives, both components attain competitive results while UNO significantly outperforms both of them.
-In the case of synthetic negatives, uncertainty outperforms negative objectness by a wide margin.
+In the case of synthetic negatives, uncertainty outperforms the negative objectness by a wide margin.
 This does not come as a surprise since the negative objectness is inferred from synthetic training negatives.
-Still, UNO greatly benefits from the ensemble of the two scores.
+Still, UNO again greatly benefits from the ensemble of the two scores.
 Score FS L&F FS Static RoadAnomaly FS L&F FS Static RoadAnomaly
@@
 Bottom row shows anomaly detection maps after thresholding at TPR=95%.
-UNO significantly reduces false-positive responses.
+UNO significantly reduces the incidence of false-positive responses.
 6 Conclusion
@@
 The resulting UNO score manifests as a lightweight plug-in module that can extend an arbitrary classifier.
-The resulting model is then fine-tuned with real or synthetic negative data that imitate test outliers.
+The resulting model is then fine-tuned with real negative data or their synthetic surrogates.
 UNO can be applied to per-pixel outlier detection by extending recent mask-level recognition approaches.
@@
 Prominent future developments include improving the quality of synthetic negative data and extending our ensemble with generative predictions.
-Limitations of UNO are discussed in the Appendix.
+7 Limitations
+Auxiliary training data.
+UNO relies on negative training data to deliver strong OOD detection performance.
+While standard applications offer plenty of real negative data, this might not be the case in domain-specific applications such as remote sensing and medical imaging.
+Still, we show that UNO can deliver strong performance when trained on synthetic negatives.
+Confidence calibration.
+Our K+1-way classifier could be inadequately calibrated which would introduce bias into our uncertainty estimates.
+An interesting direction for future work is formulation of a better uncertainty estimate.
 Acknowledgements
-This work has been co-funded by European Union - NextGenerationEU and Croatian Science Foundation grant IP-2020-02-5851 ADEPT.
+This work has been supported by Croatian Recovery and Resilience Fund - NextGenerationEU (grant C1.4 R5-I2.01.0001), and Croatian Science Foundation (grant IP-2020-02-5851 ADEPT)
+Appendix A Correlation analysis of UNO components
+Table 5 reports the Pearson correlation coefficient of per-pixel outlier scores $\textbf{s}_{\text{Unc}}$ and $\textbf{s}_{\text{NO}}$ on Fishyscapes [Blum et al.(2021)Blum, Sarlin, Nieto, Siegwart, and Cadena] val and RoadAnomaly [Lis et al.(2019)Lis, Nakka, Fua, and Salzmann].
+We observe that the two UNO components are either mildly correlated or completely uncorrelated.
+These findings indicate that the performance gains of UNO can be explained by the ensemble learning [Kuncheva and Whitaker(2003)].
+Data FS L&F FS Static RoadAnomaly
+Outliers 0.27 0.01 0.13
+Inliers 0.01 0.15 0.01
+All 0.16 0.56 0.41
+Table 5: Pearson correlation coefficient of per-pixel scores $\textbf{s}_{\text{Unc}}$ and $\textbf{s}_{\text{NO}}$ on the three outlier segmentation validation datasets (FS L&F, FS Static and RoadAnomaly).
+The benefits of ensembling can also be observed in the feature space.
+Figure 6 indicates that features from different outlier datasets have different $L_{2}$ norms.
+Outliers that resemble the training negatives typically have a higher norm and small angle to $\mathbf{w}_{K+1}$, while outliers that are more similar to inliers have lower norm.
+In the former case, outliers are detected with $\textbf{s}_{\text{NO}}$ and with $\textbf{s}_{\text{Unc}}$ in the latter case.
+Figure 6: Visualization of the pre-logit space for OpenOOD CIFAR-10.
+Outlier feature representations either yield an above-average norm with a small angle to $\mathbf{w}_{K+1}$ (e.g.
+SVHN and Places365) or yield low norm representations (e.g.
+CIFAR-100 that is similar to inliers).
+Appendix B On orthogonality of class vectors
+We empirically observe that all class vectors $\mathbf{w}_{i}$ are mutually orthogonal to each other, $\mathbf{w}_{i}^{T}\mathbf{w}_{j}=0,\,\,\forall i,j\in\{1,2,...,K+1\}$, as shown in Figure 7.
+The cosine of the angle between any two different class vectors is approximately zero.
+Such behaviour allows the geometrical interpretation of UNO, as shown in Figure 2 of the main manuscript.
+Figure 7: We show that all weight vectors are mutually orthogonal in the form of a heatmap where each element shows the cosine of the angle between two corresponding weight vectors.
+We use the K+1-way image-wide classifier with the ResNet-18 backbone trained on CIFAR-10 as the example.
+Appendix C Scene parsing with mask-wide recognition
+We extend the Mask2Former [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar] architecture with our UNO outlier detector to solve the task of anomaly segmentation.
+Thus, we describe Mask2Former architecture in detail to make the manuscript self-contained.
+The Mask2Former architecture consists of three main parts: backbone, pixel decoder, and mask decoder.
+The backbone extracts features at multiple scales from a given image $\mathbf{x}\in\mathbb{R}^{3\times H\times W}$.
+The pixel decoder produces high-resolution per-pixel features $\mathbf{E}\in\mathbb{R}^{E\times H\times W}$ that are fed into the mask decoder.
+The mask decoder formulates semantic segmentation as a direct set prediction problem by providing two outputs: $N$ mask embeddings $\mathbf{q}$, and $N$ mask-wide categorical distributions over K+1 classes $P(Y=k|\mathbf{z}_{i})=\mathrm{softmax}(\mathbf{W}\cdot\mathbf{z}_{i}+\textbf{b}).$ The K+1 classes include K inlier classes and one no-object class.
+Note that $\mathbf{z}_{i}$ denotes the vector of mask-wide pre-logit activations of the i-th mask.
+The mask decoder projects pre-logits $\mathbf{z}_{i}$ into logits by applying the learned matrix $\mathbf{W}$.
+The binary masks $\mathbf{m}=\sigma(\text{conv}_{1\times 1}(\mathbf{E},\mathbf{q}))$ are obtained by scoring per-pixel features $\mathbf{E}$ with the mask embeddings $\mathbf{q}$.
+The sigmoid activation interprets each element of the obtained tensor as a probabilistic assignment of the particular pixel into the corresponding mask.
+Semantic segmentation can be carried out by classifying each pixel according to a weighted ensemble of per-mask classifiers $P(y|\mathbf{z})$, where the weights correspond to dense mask assignments $\mathbf{m}$:
+$\hat{y}[r,c]=\underset{k=1,\dots,K}{\text{argmax}}\sum_{i}^{N}\mathbf{m}_{i}[r,c]\,P(Y=k|\mathbf{z}_{i}).$ (7)
+Default mask-level posterior already includes K+1 classes that correspond to K inlier classes and one no-object class.
+We implement our method by introducing an additional class to learn the negative objectness, which brings us to K+2 classes in total.
+To be consistent with the image-wide setup, in addition to the K inlier classes, we place the outlier class at index K+1 and the no-object class at index K+2.
+We expose the segmentation model to negative data by training on mixed-content images [Chan et al.(2021b)Chan, Rottmann, and Gottschalk].
+Closed-set recognition can still be carried out by considering only the K inlier logits (7).
+We detect anomalies on the mask-level by applying UNO to the mask-wide pre-logits $\textbf{z}_{i}$ of the i-th mask.
+We define the per-pixel outlier score at spatial positions $r$ and $c$ as a sum of mask-level outlier scores weighted with with dense probabilistic mask assignments [Grcić et al.(2023)Grcić, Šarić, and Šegvić]:
+$\textbf{s}_{\text{UNO}}^{\text{M2F}}[r,c]=\sum_{i=1}^{N}\mathbf{m}_{i}[r,c]\cdot\textbf{s}_{\text{UNO}}(\textbf{z}_{i})\;.$ (8)
+Appendix D Experimental setup
+D.1 Benchmarks and datasets
+Pixel-level benchmarks.
+Fishyscapes [Blum et al.(2021)Blum, Sarlin, Nieto, Siegwart, and Cadena] contains datasets with real (FS Lost&Found) and synthetic (FS Static) outliers.
+SMIYC [Chan et al.(2021a)Chan, Lis, Uhlemeyer, Blum, Honari, Siegwart, Fua, Salzmann, and Rottmann] has two dominant tracks which group anomalies according to size.
+AnomalyTrack focuses on the detection of large anomalies on the traffic scenes while ObstacleTrack focuses on the detection of small obstacles on the road.
+Additionaly, we validate on the RoadAnomaly [Lis et al.(2019)Lis, Nakka, Fua, and Salzmann] dataset which is an early version of the AnomalyTrack dataset.
+Image-level benchmarks.
+OpenOOD [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al.] proposed a unified benchmark for image-wide OOD detection with large-scale datasets.
+The test outliers are divided into two groups (Near-OOD and Far-OOD) based on semantical similarity to the inlier classes or observed empirical difficulty.
+The far-OOD group consists of outliers that are semantically far from the inliers (numerical digits, textural patterns, or scene imagery).
+The near-OOD group consists of outliers that are semantically similar to the inliers as they all include specific objects.
+The OpenOOD-CIFAR-10 setup uses the official CIFAR-10 [Krizhevsky(2009)] splits as ID train, val and test subsets.
+The negative dataset corresponds to a subset of Tiny ImageNet (TIN) [Le and Yang(2015)].
+The near-OOD datasets include CIFAR-100 [Krizhevsky et al.(2009)Krizhevsky, Nair, and Hinton] and a subset of Tiny ImageNet (TIN) [Le and Yang(2015)] that does not overlap with CIFAR-10 and the negative dataset.
+The far-OOD group consists of MNIST [Deng(2012)], SVHN [Yuval(2011)], Textures [Cimpoi et al.(2014)Cimpoi, Maji, Kokkinos, Mohamed, and Vedaldi], and Places365 [Zhou et al.(2017a)Zhou, Lapedriza, Khosla, Oliva, and Torralba] without images that are related with any of the ID classes.
+The large-scale OpenOOD ImageNet-200 benchmark considers a subset of 200 classes from ImageNet-1K[Deng et al.(2009)Deng, Dong, Socher, Li, Li, and Fei-Fei] as the inlier training dataset.
+The remaining 800 classes are used as the negative dataset.
+The near-OOD group consists of SSB-hard [Vaze et al.(2021)Vaze, Han, Vedaldi, and Zisserman] and NINCO [Bitterwolf et al.(2023)Bitterwolf, Müller, and Hein] while the far-OOD group includes iNaturalist [Van Horn et al.(2018)Van Horn, Mac Aodha, Song, Cui, Sun, Shepard, Adam, Perona, and Belongie], Textures [Cimpoi et al.(2014)Cimpoi, Maji, Kokkinos, Mohamed, and Vedaldi], and OpenImage-O [Wang et al.(2022)Wang, Li, Feng, and Zhang].
+D.2 Evaluation metrics
+We use standard evaluation metrics: area under the precision-recall curve (AP), area under the receiver operating curve (AUROC or AUC), and false positive rate at $95\%$ true positive rate ($\text{FPR}_{95}$).
+We validate in-distribution performance with accuracy and mIoU.
+Note that we omit the AUROC metric in the pixel-level experiments since all methods achieve high AUROC within variance.
+D.3 Implementation details
+Segmentation of road scenes.
+Our pixel-level experiments build upon a Mask2Former model [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar] with an ImageNet-initialized SWIN-L [Liu et al.(2021)Liu, Lin, Cao, Hu, Wei, Zhang, Lin, and Guo] backbone.
+We pre-train the Mask2Former in closed-set setup for 115K iterations on Cityscapes [Cordts et al.(2016)Cordts, Omran, Ramos, Rehfeld, Enzweiler, Benenson, Franke, Roth, and Schiele] and Mapillary Vistas [Neuhold et al.(2017)Neuhold, Ollmann, Rota Bulo, and Kontschieder] with the Cityscapes taxonomy.
+We use the default hyperparameters [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar] and set the batch size to 18.
+We extend the mask-wide classifier to K+2 classes and fine-tune the model on mixed-content scenes with either real or synthetic negatives for 2K iterations.
+We use the zero-initialization for the added weights.
+When training with real negative data, we assemble mixed-content images by pasting three semantically different instances sampled from ADE20K [Zhou et al.(2019)Zhou, Zhao, Puig, Xiao, Fidler, Barriuso, and Torralba] after resizing to the range of [96, 512].
+In the case of training with synthetic negatives, we jointly fine tune the K+2-way segmentation model and a flow by adding the loss defined in Equation 6 in the main manuscript to the standard optimization process [Cheng et al.(2022)Cheng, Misra, Schwing, Kirillov, and Girdhar].
+We use the DenseFlow-25-6 [Grcić et al.(2021c)Grcić, Grubišić, and Šegvić] that we pretrain on Mapillary Vistas for 300 epochs.
+We generate rectangular patches with spatial dimensions in the range of [48, 512] by sampling the jointly trained flow.
+We set the loss modulation parameter from Equation 6 to 0.03.
+We find that training our segmentation model with negatives from scratch, besides from beeing computationally expensive, leads to overfitting to negatives.
+Thus, we only finetune the model trained in closed-set manner with both real and synthetic negatives.
+Additionally, in this case we do not observe feature collapse when utilizing the joint loss (Equation 6 in the main manuscript) since we only jointly train for a small number of iterations.
+The closed-set training of Mask2Former lasts 48 hours, while the fine-tuning stage takes only 30 minutes on three A6000 GPUs.
+Image classification Our image-wide experiments follow the official training setup [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al.].
+When training with real negatives, we train the ResNet-18 [He et al.(2016)He, Zhang, Ren, and Sun] backbone with a K+1-way classification layer for 100 epochs from random initialization.
+We use the SGD optimizer with a momentum of 0.9 and a learning rate of 0.1 with cosine annealing decay schedule.
+We apply the weight decay of 0.0005.
+We set the batch size to 128 for CIFAR-10 and 256 for ImageNet-200.
+Each minibatch contains the equal ratio of all K+1 classes.
+Specifically, for CIFAR-10 the minibatch contains 117 inlier images and 11 negative images, and 255 inliers and only one negative sample for the ImageNet-200.
+When training with synthetic negatives we follow the two step procedure as explained in Section 4 of the main manuscript.
+We use the DenseFlow-25-6 [Grcić et al.(2021a)Grcić, Bevandić, and Šegvić] to generate synthetic samples.
+We pretrain the flow on inlier images, e.g.
+CIFAR-10 or ImageNet-200, for 300 epoch following the hyperparameters from [Grcić et al.(2021a)Grcić, Bevandić, and Šegvić].
+In the first step, we jointly train the flow pretrained on inlier images and a randomly initialized K-way classifier with the ResNet-18 backbone according to Equation 6 from the main manuscript.
+We train for 100 epochs and use the same hyperparameters as described above.
+In the second step, we freeze the flow and add the K+1-th logit to the classifier and finetune for 20 epochs.
+We set the learning rate for the backbone to 0.0001 and 0.01 for the final fully connected layer.
+We set the loss modulation parameter from Equation 6 to 0.03.
+Appendix E Additional results
+We provide a discussion on the choice of negative data, an alternative implementation of negative objectness and the full results on the OpenOOD [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al.] benchmark.
+E.1 Impact of synthetic negatives
+Table 6 shows the performance of UNO depending on the source of negative training data.
+We experiment with random crops from inlier scenes, synthetic negatives generated by a jointly trained normalizing flow, and random instances from the ADE datasets.
+Training on synthetic negatives is more beneficial than training on inlier even though the flow was jointly trained only for a small number of iterations.
+For instance, training on inlier crops yields a high FPR95 on Fishyscapes Lost&Found.
+Contrary, synthetic negatives yield low FPR95 on all three validations sets.
+Still, there is a performance gap between models that are trained with and without real negative data.
+Table 6: Performance of UNO for different training negatives in per-pixel outlier segmentation.
+Training FS L&F FS Static RoadAnomaly
+negatives AP $\text{FPR}_{95}$ AP $\text{FPR}_{95}$ AP $\text{FPR}_{95}$
+Inlier crops 67.2 62.5 79.2 0.9 86.5 7.6
+Synthetic negatives 74.5 6.9 96.9 0.1 82.4 9.2
+ADE20k negatives 81.8 1.3 98.0 0.04 88.5 7.4
+E.2 Alternative implementation of the outlier posterior
+The outlier posterior $P(y_{\text{NO}}|\mathbf{z})$ can alternatively be modeled with an additional out-of-distribution head [Bevandić et al.(2022)Bevandić, Krešo, Oršić, and Šegvić].
+Then, we introduce an additional binary cross-entropy loss term to train the out-of-distribution head that discriminates inliers and outliers.
+This way the closed set classifier ends up with K classes and is not affected by the negative data.
+When applied to the mask-wide recognition architecture, the OOD head has 3 outputs 1) inlier, 2) outlier and 3) no-object.
+Table 7 shows the comparison of the K+2-way classifier proposed in the main paper with a K+1-way classifier and a 3-way OOD head and ablation of UNO components in the pixel-wise outlier detection setup.
+Our original UNO formulation built atop K+2-way classifier consistently outperforms alternative formulations across all datasets and metrics.
+Table 7: Comparison of the K+2-way classifier with the OOD head atop of Mask2Former architecture on Fishyscapes val and RoadAnomaly.
+FS L&F FS Static RoadAnomaly
+Method Score AP $\text{FPR}_{95}$ AP $\text{FPR}_{95}$ AP $\text{FPR}_{95}$
+K+2-way classifier $\text{s}_{\text{UNO}}$ 81.8 1.3 98.0 0.0 88.5 7.4
+K-way classifier & OOD head $\text{s}_{\text{UNO}}$ 81.4 5.3 89.0 0.3 80.6 10.6
+K-way classifier & OOD head $\text{s}_{\text{Unc}}$ 77.8 2.2 86.4 1.6 76.1 8.4
+K-way classifier & OOD head $\text{s}_{\text{NO}}$ 79.9 7.3 88.3 0.3 74.5 25.2
+Table 8 presents a similar analysis in the image-wide setting using the OpenOOD CIFAR-10 setup.
+Again, the K+1-way classifier combined with our UNO score outperforms the K-way classifier and a binary OOD head.
+Still, our UNO score works well even with the alternative formulation.
+Table 8: Comparison of the K+1-way classifier with the binary OOD head atop of ResNet-18 on OpenOOD CIFAR-10. We use UNO as the outlier score.
+Method Near-OOD Far-OOD
+AUC $\text{FPR}_{95}$ AUC $\text{FPR}_{95}$
+K+1-way classifier 95.00 18.75 97.95 9.30
+K-way classifier & OOD head 94.23 19.31 97.76 11.12
+E.3 Full results on OpenOOD
+Tables 9 and 10 provide the extended results on the OpenOOD [Zhang et al.(2023b)Zhang, Yang, Wang, Wang, Lin, Zhang, Sun, Du, Zhou, Zhang, et al.] benchmark.
+We compare with post-hoc methods (upper section) and training methods without (middle section) and with the use of real negative data (bottom section).
+All results are averaged over three runs with variances in subscripts.
+Table 9: OOD detection performance on the OpenOOD benchmark, the CIFAR-10 dataset.
+Method Near-OOD Far-OOD
+AUC FPR95 AP AUC FPR95 AP Acc.
+OpenMax [Bendale and Boult(2016)] 87.62(±0.29) 43.62(±2.27) 80.60(±0.29) 89.62(±0.19) 29.69(±1.21) 90.19(±0.41) 95.06(±0.30)
+MSP [Hendrycks and Gimpel(2017)] 88.03(±0.25) 48.17(±3.92) 85.43(±0.36) 90.73(±0.43) 31.72(±1.84) 93.27(±0.14) 95.06(±0.30)
+TempScale [Guo et al.(2017)Guo, Pleiss, Sun, and Weinberger] 88.09(±0.31) 50.96(±4.32) 86.11(±0.33) 90.97(±0.52) 33.48(±2.39) 93.68(±0.11) 95.06(±0.30)
+ODIN [Liang et al.(2017)Liang, Li, and Srikant] 82.87(±1.85) 76.19(±6.08) 83.03(±1.15) 87.96(±0.61) 57.62(±4.24) 93.14(±0.44) 95.06(±0.30)
+MDS [Lee et al.(2018b)Lee, Lee, Lee, and Shin] 84.20(±2.40) 49.90(±3.98) 79.88(±3.18) 89.72(±1.36) 32.22(±3.40) 93.81(±0.74) 95.06(±0.30)
+MDSEns [Lee et al.(2018b)Lee, Lee, Lee, and Shin] 60.43(±0.26) 92.26(±0.20) 59.94(±0.19) 73.90(±0.27) 61.47(±0.48) 83.37(±0.04) 95.06(±0.30)
+RMDS [Ren et al.(2021)Ren, Fort, Liu, Roy, Padhy, and Lakshminarayanan] 89.80(±0.28) 38.89(±2.39) 87.52(±0.29) 92.20(±0.21) 25.35(±0.73) 94.21(±0.10) 95.06(±0.30)
+Gram [Sastry and Oore(2020)] 58.66(±4.83) 90.87(±1.91) 57.57(±5.09) 71.73(±3.20) 72.34(±6.73) 82.89(±3.14) 95.06(±0.30)
+EBO [Liu et al.(2020)Liu, Wang, Owens, and Li] 87.58(±0.46) 61.34(±4.63) 87.04(±0.27) 91.21(±0.92) 41.69(±5.32) 94.31(±0.09) 95.06(±0.30)
+OpenGAN [Kong and Ramanan(2021)] 53.71(±7.68) 94.48(±4.01) 53.35(±5.22) 54.61(±15.51) 83.52(±11.63) 73.34(±8.49) 95.06(±0.30)
+GradNorm [Huang et al.(2021)Huang, Geng, and Li] 54.90(±0.98) 94.72(±0.82) 57.95(±1.98) 57.55(±3.22) 91.90(±2.23) 76.75(±1.95) 95.06(±0.30)
+ReAct [Sun et al.(2021)Sun, Guo, and Li] 87.11(±0.61) 63.56(±7.33) 86.65(±0.19) 90.42(±1.41) 44.90(±8.37) 93.99(±0.45) 95.06(±0.30)
+MLS [Hendrycks et al.(2022b)Hendrycks, Basart, Mazeika, Zou, Kwon, Mostajabi, Steinhardt, and Song] 87.52(±0.47) 61.32(±4.62) 86.88(±0.29) 91.10(±0.89) 41.68(±5.27) 94.21(±0.06) 95.06(±0.30)
+KLM [Hendrycks et al.(2022b)Hendrycks, Basart, Mazeika, Zou, Kwon, Mostajabi, Steinhardt, and Song] 79.19(±0.80) 87.86(±6.37) 80.37(±0.46) 82.68(±0.21) 78.31(±4.84) 90.57(±0.25) 95.06(±0.30)
+VIM [Wang et al.(2022)Wang, Li, Feng, and Zhang] 88.68(±0.28) 44.84(±2.31) 86.32(±0.39) 93.48(±0.24) 25.05(±0.52) 96.27(±0.24) 95.06(±0.30)
+KNN [Sun et al.(2022)Sun, Ming, Zhu, and Li] 90.64(±0.20) 34.01(±0.38) 88.50(±0.35) 92.96(±0.14) 24.27(±0.40) 94.93(±0.07) 95.06(±0.30)
+DICE [Sun and Li(2022)] 78.34(±0.79) 70.04(±7.64) 74.80(±2.33) 84.23(±1.89) 51.76(±4.42) 89.06(±1.72) 95.06(±0.30)
+RankFeat [Song et al.(2022)Song, Sebe, and Wang] 79.46(±2.52) 60.88(±4.60) 74.46(±3.08) 75.87(±5.06) 57.44(±7.99) 81.27(±3.67) 95.06(±0.30)
+ASH [Djurisic et al.(2022)Djurisic, Bozanic, Ashok, and Liu] 75.27(±1.04) 86.78(±1.82) 77.24(±1.26) 78.49(±2.58) 79.03(±4.22) 88.33(±1.39) 95.06(±0.30)
+SHE [Zhang et al.(2023c)Zhang, Fu, Chen, Du, Li, Wang, Liu, Han, and Zhang] 81.54(±0.51) 79.65(±3.47) 82.04(±0.51) 85.32(±1.43) 66.48(±5.98) 91.26(±0.04) 95.06(±0.30)
+ConfBranch [DeVries and Taylor(2018)] 89.84(±0.24) 31.28(±0.66) 85.50(±0.30) 92.85(±0.29) 94.88(±0.05) 93.48(±0.39) 94.88(±0.05)
+RotPred [Hendrycks et al.(2019b)Hendrycks, Mazeika, Kadavath, and Song] 92.68(±0.27) 28.14(±1.68) 90.47(±0.35) 96.62(±0.18) 12.23(±0.33) 97.54(±0.13) 95.35(±0.52)
+G-ODIN [Hsu et al.(2020)Hsu, Shen, Jin, and Kira] 89.12(±0.57) 45.54(±2.52) 88.25(±0.49) 95.51(±0.31) 21.45(±1.91) 97.35(±0.34) 94.70(±0.25)
+CSI [Tack et al.(2020)Tack, Mo, Jeong, and Shin] 89.51(±0.19) 33.66(±0.64) 86.37(±0.25) 92.00(±0.30) 26.42(±0.29) 93.90(±0.33) 91.16(±0.14)
+ARPL [Chen et al.(2022)Chen, Peng, Wang, and Tian] 87.44(±0.15) 40.33(±0.70) 82.96(±0.33) 89.31(±0.32) 32.39(±0.74) 91.41(±0.09) 93.66(±0.11)
+MOS [Huang and Li(2021)] 71.45(±3.09) 78.72(±5.86) 72.41(±3.05) 76.41(±5.93) 62.90(±6.62) 85.24(±2.92) 94.83(±0.37)
+VOS [Du et al.(2022)Du, Wang, Cai, and Li] 87.70(±0.48) 57.03(±1.92) 86.57(±0.73) 90.83(±0.92) 40.43(±4.53) 93.95(±0.56) 94.31(±0.64)
+LogitNorm [Wei et al.(2022)Wei, Xie, Cheng, Feng, An, and Li] 92.33(±0.08) 29.34(±0.81) 90.62(±0.09) 96.74(±0.06) 13.81(±0.20) 97.29(±0.21) 94.30(±0.25)
+CIDER [Ming et al.(2023)Ming, Sun, Dia, and Li] 90.71(±0.16) 32.11(±0.94) 87.97(±0.24) 94.71(±0.36) 20.72(±0.85) 96.19(±0.19) -
+NPOS [Tao et al.(2023)Tao, Du, Zhu, and Li] 89.78(±0.33) 32.64(±0.70) 86.36(±0.68) 94.07(±0.49) 20.59(±0.69) 96.20(±0.43) -
+UNO (ours) 91.34(±0.33) 31.78(±0.82) 87.39(±0.35) 92.55(±0.25) 20.54(±0.87) 93.96(±0.25) 95.20(±0.25)
+MixOE [Zhang et al.(2023a)Zhang, Inkawhich, Linderman, Chen, and Li] 88.73(±0.82) 51.45(±7.78) 94.25(±0.17) 91.93(±0.69) 33.84(±4.77) 98.00(±0.02) 94.55(±0.32)
+MCD [Yu and Aizawa(2019)] 91.03(±0.12) 30.17(±0.06) 87.73(±0.36) 91.00(±1.10) 32.03(±4.21) 94.45(±0.85) 94.95(±0.04)
+UDG [Yang et al.(2021)Yang, Wang, Feng, Yan, Zheng, Zhang, and Liu] 89.91(±0.25) 35.34(±0.95) 86.89(±0.84) 94.06(±0.90) 20.35(±2.41) 95.23(±0.89) 92.36(±0.84)
+OE [Hendrycks and Gimpel(2017)] 94.82(±0.21) 19.84(±0.95) 87.39(±0.60) 96.00(±0.13) 13.13(±0.53) 95.03(±0.12) 94.63(±0.26)
+UNO (ours) 94.87(±0.07) 9.33(±0.50) 94.49(±0.13) 97.63(±0.72) 9.38(±2.65) 99.10(±0.25) 94.88(±0.19)
+Table 10: OOD detection performace on the OpenOOD benchmark, the Imagenet-200 dataset.
+Method Near-OOD Far-OOD
+AUC FPR95 AP AUC FPR95 AP Acc.
+OpenMax [Bendale and Boult(2016)] 80.27(±0.10) 63.48(±0.25) 81.42(±0.19) 90.20(±0.17) 33.12(±0.66) 85.13(±0.47) 86.37(±0.08)
+MSP [Hendrycks and Gimpel(2017)] 83.34(±0.06) 54.82(±0.35) 85.95(±0.05) 90.13(±0.09) 35.43(±0.38) 88.71(±0.14) 86.37(±0.08)
+TempScale [Guo et al.(2017)Guo, Pleiss, Sun, and Weinberger] 83.69(±0.04) 54.82(±0.23) 86.29(±0.02) 90.82(±0.09) 34.00(±0.37) 89.49(±0.15) 86.37(±0.08)
+ODIN [Liang et al.(2017)Liang, Li, and Srikant] 80.27(±0.08) 66.76(±0.26) 85.02(±0.03) 91.71(±0.19) 34.23(±1.05) 91.29(±0.15) 86.37(±0.08)
+MDS [Lee et al.(2018b)Lee, Lee, Lee, and Shin] 61.93(±0.51) 79.11(±0.31) 67.68(±0.42) 74.72(±0.26) 61.66(±0.27) 70.80(±0.61) 86.37(±0.08)
+MDSEns [Lee et al.(2018b)Lee, Lee, Lee, and Shin] 54.32(±0.24) 91.75(±0.10) 64.81(±0.24) 69.27(±0.57) 80.96(±0.38) 69.62(±0.52) 86.37(±0.08)
+RMDS [Ren et al.(2021)Ren, Fort, Liu, Roy, Padhy, and Lakshminarayanan] 82.57(±0.25) 54.02(±0.58) 83.07(±0.45) 88.06(±0.34) 32.45(±0.79) 82.71(±0.78) 86.37(±0.08)
+Gram [Sastry and Oore(2020)] 67.67(±1.07) 86.40(±1.21) 75.63(±0.78) 71.19(±0.24) 84.36(±0.78) 72.75(±0.25) 86.37(±0.08)
+EBO [Liu et al.(2020)Liu, Wang, Owens, and Li] 82.50(±0.05) 60.24(±0.57) 85.48(±0.07) 90.86(±0.21) 34.86(±1.30) 89.85(±0.21) 86.37(±0.08)
+OpenGAN [Kong and Ramanan(2021)] 59.79(±3.39) 84.15(±3.85) 66.85(±2.79) 73.15(±4.07) 64.16(±9.33) 66.62(±3.69) 86.37(±0.08)
+GradNorm [Huang et al.(2021)Huang, Geng, and Li] 72.75(±0.48) 82.67(±0.30) 80.19(±0.68) 84.26(±0.87) 66.45(±0.22) 86.54(±0.92) 86.37(±0.08)
+ReAct [Sun et al.(2021)Sun, Guo, and Li] 81.87(±0.98) 62.49(±2.19) 85.38(±0.34) 92.31(±0.56) 28.50(±0.95) 91.31(±0.80) 86.37(±0.08)
+MLS [Hendrycks et al.(2022b)Hendrycks, Basart, Mazeika, Zou, Kwon, Mostajabi, Steinhardt, and Song] 82.90(±0.04) 59.76(±0.59) 85.96(±0.07) 91.11(±0.19) 34.03(±1.21) 90.10(±0.21) 86.37(±0.08)
+KLM [Hendrycks et al.(2022b)Hendrycks, Basart, Mazeika, Zou, Kwon, Mostajabi, Steinhardt, and Song] 80.76(±0.08) 70.26(±0.64) 83.41(±0.23) 88.53(±0.11) 40.90(±1.08) 84.22(±0.47) 86.37(±0.08)
+VIM [Wang et al.(2022)Wang, Li, Feng, and Zhang] 78.68(±0.24) 59.19(±0.71) 81.61(±0.29) 91.26(±0.19) 27.20(±0.30) 90.01(±0.35) 86.37(±0.08)
+KNN [Sun et al.(2022)Sun, Ming, Zhu, and Li] 81.57(±0.17) 60.18(±0.52) 85.72(±0.17) 93.16(±0.22) 27.27(±0.75) 93.48(±0.15) 86.37(±0.08)
+DICE [Sun and Li(2022)] 81.78(±0.14) 61.88(±0.67) 85.37(±0.13) 90.80(±0.31) 36.51(±1.18) 90.55(±0.29) 86.37(±0.08)
+RankFeat [Song et al.(2022)Song, Sebe, and Wang] 56.92(±1.59) 92.06(±0.23) 66.17(±1.63) 38.22(±3.85) 97.72(±0.75) 45.25(±2.81) 86.37(±0.08)
+ASH [Djurisic et al.(2022)Djurisic, Bozanic, Ashok, and Liu] 82.38(±0.19) 64.89(±0.90) 87.03(±0.06) 93.90(±0.27) 27.29(±1.12) 94.15(±0.32) 86.37(±0.08)
+SHE [Zhang et al.(2023c)Zhang, Fu, Chen, Du, Li, Wang, Liu, Han, and Zhang] 80.18(±0.25) 66.80(±0.74) 84.20(±0.28) 89.81(±0.61) 42.17(±1.24) 90.05(±0.62) 86.37(±0.08)
+ConfBranch [DeVries and Taylor(2018)] 79.10(±0.24) 61.44(±0.34) 82.11(±0.30) 90.43(±0.18) 34.75(±0.63) 88.67(±0.27) 85.92(±0.07)
+RotPred [Hendrycks et al.(2019b)Hendrycks, Mazeika, Kadavath, and Song] 81.59(±0.20) 60.42(±0.60) 84.87(±0.19) 92.56(±0.09) 26.16(±0.38) 90.10(±0.08) 86.37(±0.16)
+G-ODIN [Hsu et al.(2020)Hsu, Shen, Jin, and Kira] 77.28(±0.10) 69.87(±0.46) 82.77(±0.16) 92.33(±0.11) 30.18(±0.49) 92.04(±0.10) 84.56(±0.28)
+ARPL [Chen et al.(2022)Chen, Peng, Wang, and Tian] 82.02(±0.10) 55.74(±0.70) 84.35(±0.08) 89.23(±0.11) 36.46(±0.08) 87.63(±0.19) 83.95(±0.32)
+MOS [Huang and Li(2021)] 69.84(±0.46) 71.60(±0.48) 73.38(±0.56) 80.46(±0.92) 51.56(±0.42) 72.79(±1.43) 85.60(±0.20)
+VOS [Du et al.(2022)Du, Wang, Cai, and Li] 82.51(±0.11) 59.89(±0.47) 85.59(±0.07) 91.00(±0.28) 34.01(±0.97) 90.11(±0.30) 86.23(±0.19)
+LogitNorm [Wei et al.(2022)Wei, Xie, Cheng, Feng, An, and Li] 82.66(±0.15) 56.46(±0.37) 86.41(±0.08) 93.04(±0.21) 26.11(±0.52) 92.25(±0.32) 86.04(±0.15)
+CIDER [Ming et al.(2023)Ming, Sun, Dia, and Li] 80.58(±1.75) 60.10(±0.73) 83.32(±1.76) 90.66(±1.68) 30.17(±2.75) 89.16(±2.38) -
+NPOS [Tao et al.(2023)Tao, Du, Zhu, and Li] 79.40(±0.39) 62.09(±0.05) 84.37(±0.35) 94.49(±0.07) 21.76(±0.21) 94.83(±0.07) -
+UNO (ours) 81.16(±0.65) 61.10(±0.93) 85.31(±0.94) 92.53(±0.48) 32.32(±0.19) 87.24(±0.33) 86.33(±0.36)
+OE [Hendrycks and Gimpel(2017)] 84.84(±0.16) 52.30(±0.67) 86.86(±0.22) 89.02(±0.18) 34.17(±0.56) 85.15(±0.26) 85.82(±0.21)
+MCD [Yu and Aizawa(2019)] 83.62(±0.09) 54.71(±0.83) 84.44(±0.30) 88.94(±0.10) 29.93(±0.30) 82.90(±0.48) 86.12(±0.17)
+UDG [Yang et al.(2021)Yang, Wang, Feng, Yan, Zheng, Zhang, and Liu] 74.30(±1.63) 68.89(±1.72) 78.09(±2.02) 82.09(±2.78) 62.04(±5.99) 81.63(±3.29) 68.11(±1.24)
+MixOE [Zhang et al.(2023a)Zhang, Inkawhich, Linderman, Chen, and Li] 82.62(±0.03) 57.97(±0.40) 84.78(±0.05) 88.27(±0.41) 40.93(±0.29) 86.01(±0.60) 85.71(±0.07)
+UNO (ours) 85.07(±0.78) 51.71(±0.31) 85.29(±0.71) 89.63(±0.46) 36.79(±0.21) 87.07(±0.14) 86.42(±0.32)